### v2

# Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection

Abstract
The spread of fake news negatively impacts individuals and is regarded as a significant social challenge that needs to be addressed. A number of algorithmic and insightful features have been identified for detecting fake news. However, with the recent LLMs and their advanced generation capabilities, many of the detectable features (e.g., style-conversion attacks) can be altered, making it more challenging to distinguish from real news. This study proposes adversarial style augmentation, AdStyle, to train a fake news detector that remains robust against various style-conversion attacks. Our model’s key mechanism is the careful use of LLMs to automatically generate a diverse yet coherent range of style-conversion attack prompts. This improves the generation of prompts that are particularly difficult for the detector to handle. Experiments show that our augmentation strategy improves robustness and detection performance when tested on fake news benchmark datasets.
Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection
Sungwon Park1, Sungwon Han1, Meeyoung Cha2 1School of Computing, KAIST, Daejeon, Republic of Korea 2MPI for Security and Privacy, Bochum, Germany {psw0416, lion4151}@kaist.ac.kr, mia.cha@mpi-sp.org
1 Introduction
With the widespread use of the internet and the emergence of various social media platforms, people have begun to freely share information they know and their own creative stories. However, the ease of sharing and consuming information beyond traditional news organizations has also significantly contributed to the spread of fake news Cha et al. (2021); Fisher et al. (2016); Treen et al. (2020). Fake news is created to provide readers with false information as propaganda or to guide public perception in a desired direction for intentional objectives Lazer et al. (2018); Yi et al. (2024). The spread of such false information has had profound negative impacts on individuals and society, becoming a major social challenge that needs to be addressed.
Manually determining the authenticity of all information on the internet is extremely costly in terms of time and resources. Thus, prior advances have mainly focused on developing automated fake news detectors using machine learning techniques Pérez-Rosas et al. (2018); Reis et al. (2019); Shu et al. (2019). For example, several studies have extracted sentiment-related or political features commonly found in fake news, either through human-crafted Potthast et al. (2018); Rashkin et al. (2017) or data-driven methods Reis et al. (2019); Shu et al. (2019), and used them for detection. With the advent of language models (e.g., GPT, BERT) Kenton and Toutanova (2019); Radford et al. (2019), efforts have been made to learn the textual style of fake news by extracting sentence embeddings to train detectors Chen et al. (2021); Yang et al. (2022).
While these extracted textual styles or human-crafted features have proven effective in identifying fake news, they allow attackers to bypass detection. In particular, the advent of large language models (LLMs) Chowdhery et al. (2022); OpenAI (2023) has made it possible to easily and automatically paraphrase sentences in a user-desired direction through prompts (e.g., "change the given text to an objective and professional style"), making it more difficult to distinguish between AI-generated fake news and real news (i.e., style-conversion attacks) Koenders et al. (2021); Zhou et al. (2019). Recent literature has proposed strategies to augment the styles of input text via some manually defined style-conversion prompts, which are taken by LLM for paraphrasing, but it remains challenging to address all types of prompts that attackers might use to deceive detectors Wu and Hooi (2023).
In this study, we propose adversarial style augmentation, AdStyle, to train a robust fake news detector that can withstand various style-conversion attacks by attackers. In contrast to earlier work, which used predefined style-conversion prompts as detector-agnostic augmentations, AdStyle tries to find prompts for adversarial style augmentations that are specific to the detector, which makes predictions uncertain. This means that our augmentations add noise to the style features in the direction of the detector’s decision boundary, while maintaining the text’s content integrity. Specifically, to search for and optimize prompts for LLM that are adversarial to the detector, we introduce an automated prompt engineering technique Yang et al. (2023). By providing LLMs with style-conversion prompts and the detector’s performance under these augmentations, LLM can infer patterns between the prompts and performance, enabling the search for the most adversarial prompts.
AdStyle proceeds as follows: To generate augmented samples, we first select a random subset of the dataset and apply a pool of style-conversion prompts to it. Then, each augmentation set created by different prompts is fed into the detector, and the score based on the AUC between the predictions and the ground-truth labels is measured. The prompt-score pairs are provided to the LLM, which uses this information to generate new style-conversion prompt candidates. From these candidates, we select the top-$k$ prompts that are diverse and make the detector’s predictions most uncertain without significantly altering the original content. These selected prompts are applied to the entire dataset and used for detector training. The chosen prompts are added to the prompt pool, which is repeated over several training rounds.
We experimented with existing fake news benchmark datasets under various style-conversion attack scenarios. As a result, our augmentation strategy demonstrated higher robustness and detection performance compared to previous methodologies. Additionally, our approach is scalable to different attack strategies by adding new attack prompts to the style-conversion prompt pool. We plan to release our code after publication.
2 Related Work
2.1 Automated Detection of Fake News
Manually detecting fake news among the vast amount of content on the internet is intractable. Consequently, many studies have focused on leveraging machine learning techniques to automatically detect fake news Pérez-Rosas et al. (2018); Reis et al. (2019); Shu et al. (2019). For example, supervised learning methodologies extract textual features from fake news texts using benchmark datasets and ground-truth labels Reis et al. (2019). These textual features can include deep features based on artificial neural networks Reis et al. (2019); Shu et al. (2019) as well as manually defined features such as sentiment and political bias Potthast et al. (2018); Rashkin et al. (2017). Other approaches include domain adaptation methods to enhance detection generalizability across various domains and topics Mosallanezhad et al. (2022); Nan et al. (2022), and knowledge-based methods that rely on external data to distinguish false information Dun et al. (2021). With the advent of LLMs, new methods have emerged that utilize LLMs’ prior knowledge to identify fake news Hu et al. (2024), including research on detecting machine-generated fake news Chen and Shu (2023).
In this work, we aim to prevent malicious attackers from deceiving automated fake news detectors through text modifications such as paraphrasing. The proposed method is an augmentation strategy to enhance robustness against text style perturbations as an add-on to existing detection models. Our contribution is agnostic to the design of the detection model.
Figure 1: Illustration of how AdStyle works. Our model consists of multiple rounds of training. In each round, the model uses the style-conversion prompt and prediction confusion score from the previous round to generate prompt candidates that can maximize the confusion of the detector model. Subsequently, considering aspects such as adversarialness, coherency, and diversity, a subset of the training dataset is used to select the top-$k$ prompts, which are finally used as augmentations in training.
2.2 Attack on Fake News Detection
To test the robustness of fake news detection, various attack methods have been studied Koenders et al. (2021); Zhou et al. (2019). These include injecting misinformation by changing the order of subjects and objects or causes and effects while maintaining the textual features used for detection Koenders et al. (2021). Other methods involve creating fact distortions by altering or exaggerating words related to people, time, or places while preserving the sentence structure Zhou et al. (2019). Additionally, approaches that use the text generation capabilities of LLMs to change the style of sentences have been proposed Wu and Hooi (2023).
Our method aims to perform adversarial training of the detection model by generating augmentations that perturb the text’s style while preserving the content as much as possible. This approach reduces the impact of spurious style features on the detection model, making it more robust against such attacks.
2.3 Prompt Engineering for LLM
Training with extensive text corpora, LLMs have demonstrated their utility across various domain tasks Chowdhery et al. (2022); OpenAI (2023); Pang et al. (2024). To better harness the prior knowledge and reasoning abilities of LLMs, strategies for providing appropriate input prompts have also been researched. For instance, in-context learning methods involve providing examples of the desired input-output format to guide the LLM in producing the correct output Dong et al. (2022). Additionally, engineering techniques such as chain-of-thought prompting, which includes additional reasoning steps in the responses, have been developed to enhance the reasoning abilities of LLMs Wei et al. (2022). Recently, methods for automatically finding the most suitable prompts for a given task, known as automated prompt engineering, have shown promising results Yang et al. (2023); Zhou et al. (2022).
In this paper, we adopt automated prompt engineering techniques to identify adversarial prompts that most effectively confuse the fake news detector. This process replaces the traditional gradient descent method used in the image domain to find adversarial noise, enabling the creation of adversarially augmented versions of input text via LLM.
3 Method
3.1 Overview
Let $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$ be a dataset containing news $\mathbf{d}_{i}$ and the corresponding ground-truth binary veracity label $y_{i}$ (indicating whether the news is true or fake). Each news item $\mathbf{d}_{i}$ is composed of natural language-based text. This study aims to train a language model based fake news detector $f$ using the labeled dataset to predict the veracity labels. Our main goal is to develop a detector that remains robust even when an attacker perturbs the textual style, such as the order and format, while preserving the meaning of the sentences.
Figure 1 illustrates how our model works. AdStyle generates style-conversion prompts and performs augmentation over multiple rounds. Each round includes a following process. Firstly, leveraging the reasoning ability of LLM, adversarial style-conversion prompt candidates are generated. These style-conversion prompts contain instructions on how to transform the given sentences and are used as inputs to the LLM along with the original sentences for conversion. Inspired by automated prompt engineering Yang et al. (2023), the style-conversion prompts and detector prediction score pairs used in previous rounds are included as in-context demonstrations to guide the LLM in searching prompt candidates that maximally confuse the detector (Section 3.2). Next, a subset of the dataset is selected, and these discovered candidates are applied to perform conversions. The converted samples are then evaluated to determine how much they confuse the detector. From these candidates, the top-$k$ prompts that are diverse and maintain the original content’s meaning while most effectively confusing the detector are selected (Section 3.3). These selected prompts are used as an augmentation method to train the detection model in the current round. We describe each step’s details below.
3.2 Generating Adversarial Style-Conversion Prompts with LLM
The style-conversion prompts we aim to generate are instructions that perturb only the textual style, such as the structure or format of the sentence, while preserving the content of the given input text. For example, an instruction like “Rewrite the following article in an objective and professional tone” can be used to transform the stylistic features of fake news to resemble real news. However, if such instructions are heuristically defined, they may not align with the actual decision boundary of the detector, reducing the training efficiency. Furthermore, if the instructions remain fixed throughout the training, the detector may memorize these conversion patterns, resulting in overfitting.
Therefore, instead of predefining and fixing the instructions for conversion, we aim to find adversarial conversion prompts that add noise in the direction of the decision boundary of current state of detector, maximizing the confusion in the detector’s predictions. This approach is similar to adversarial training commonly used in the computer vision domain Madry et al. (2018), which enhances robustness against slight perturbations in the input. However, for textual data, the discrete nature of the input makes it challenging to add noise directly through the detector’s gradient as in the image domain. To address this, we measure the confusion in prediction by conversion prompts (i.e., prediction confusion score), and feed this prompt-score pair to LLM for generating adversarial prompts, motivated from automated prompt engineering techniques.
Now you will help me minimize a fake news detector performance value with style transfer attack instruction. I have some style transfer attack instructions and the fake news detector performance value using those instructions. The pairs are arranged in descending order based on their function values, where lower values are better. Input: [Style-Conversion Prompt Example] Score: [Prediction Confusion Score] Input: … … Give me a new style transfer attack instruction that is different from all pairs above, and has a performance value lower than any of the above.
Figure 2: Prompt for generating adversarial style-conversion prompts. The text in blue represents the score trajectory component, while the remaining text represents the problem description component.
Figure 2 shows an example of the LLM input used to generate adversarial style-conversion prompts. The LLM input consists of the problem description and the score trajectory components.
Problem description component.
This component includes the problem description, the objective, and constraints on the response necessary for generating style-conversion prompts. For example, a sentence like “Minimize a fake news detector performance value with style transfer attack instruction” informs the LLM of the intent behind the conversion prompt.
Score trajectory component.
Previous work has shown that LLMs can learn patterns from in-context demonstrations provided as input Dong et al. (2022); Yang et al. (2023). This component leverages this ability by providing previous round’s style-conversion prompts and their corresponding prediction confusion scores in the form of in-context demonstrations [^1] In the first round, a predefined set of prompts is used.. The score is measured by selecting a subset $\mathcal{B}$ from the entire training dataset $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$, applying a conversion prompt $c$ to create a new set $\mathcal{B}^{c}=\{(\mathbf{d}^{c}_{i},y_{i})\}_{i=1}^{M}$, where $\mathbf{d}^{c}_{i}=\text{Convert}(c,\mathbf{d}_{i}),N>>M$, and then measuring the AUC score between the predictions and the ground-truth labels when $\mathcal{B}^{c}$ is fed into the detector. Specifically, the score of a conversion prompt $s_{c}$ is defined as:
$s_{c}=|0.5-\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})|.$ (1)
A lower score indicates a higher level of prediction confusion, implying that the conversion prompt has caused the detector’s predictions to become random with respect to the labels. Finding conversion prompts that cause high confusion (i.e., low score) suggests that the detector has not yet learned to handle those stylistic features, and the conversion prompts have placed the samples near the detector’s decision boundary, making them difficult to distinguish. By showing the LLM these in-context demonstrations, it can generate conversion prompts that differ from past ones while maximizing confusion (i.e., minimizing the score). Note that we avoid selecting conversion prompts that flip the detector’s original predictions (i.e., AUC <<much-less-than$<$ < 0.5), as these often disrupt the content along with the stylistic features, which is undesirable. We extract $S$ style-conversion prompts at a time using the input described above.
3.3 Selecting Top-$k$ Adversarial Prompts
Using all the style-conversion prompt candidates generated by the LLM can be computationally intensive, and not all prompts may be suitable for augmentation. For example, the set of style-conversion prompts used for augmentation should each provide adversarial perturbations that confuse the detector (i.e., adversarialness). Additionally, while altering the textual style, the prompts should not change the meaning of the sentences to prevent label noise (i.e., coherency). Furthermore, the more diverse the set of conversion directions covered by the prompts, the more efficient the augmentation process (i.e., diversity).
To select a set of style-conversion prompts that satisfies these three criteria, we propose a selection strategy. Given a conversion prompt $c$, we first extract embedding vectors for the input texts in both the training subset $\mathcal{B}$ and the subset $\mathcal{B}^{c}$ converted by $c$ using a large language model $g$ such as BERT. We then compute the average embedding vectors for each subset and calculate the vector difference $\mathbf{z}_{c}$.
$\mathbf{z}$ $=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{d}_{i}\in\mathcal{B}}{g(\mathbf{d}_{i})},\ \ \ \mathbf{z}\textquoteright=\frac{1}{|\mathcal{B}\textquoteright|}\sum_{\mathbf{d}^{c}_{i}\in\mathcal{B}\textquoteright}{g(\mathbf{d}^{c}_{i})}$
$\mathbf{z}_{c}$ $=\mathbf{z}\textquoteright-\mathbf{z}$ (2)
Here, $\mathbf{z}_{c}$ represents the average change in embedding direction due to the conversion prompt $c$. We then calculate the adversarialness scale $s^{c}_{\text{adv}}$ and coherency scale $s^{c}_{\text{coh}}$ for each conversion prompt $c$, and rescale $\mathbf{z}_{c}$ accordingly (i.e., $\mathbf{\hat{z}}_{c}=\mathbf{z}_{c}\times s^{c}_{\text{adv}}\times s^{c}_{\text{coh}}$). Finally, we use the $k$-means++ initialization method Arthur et al. (2007) on these rescaled vectors to select $k$ prompts. The $k$-means++ initialization method helps select a diverse set of prompts that are adversarial and coherent, by choosing samples that are as far apart as possible Ash et al. (2019). The details of each scale are described below.
Adversarialness scale.
To measure how adversarial a given style-conversion prompt is to the detector, we define a new adversarialness scale similar to the confusion score defined in the previous section. Given the converted batch $\mathcal{B}^{c}$ by conversion prompt $c$, the adversarialness scale $s^{c}_{\text{adv}}$ is defined as $-1.8\cdot|\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})-0.5|+1$. This value increases as the AUC approaches 0.5, indicating that the prediction is more random. The coefficient 1.8 ensures that the scale ranges between 0.1 and 1, preventing it from being zero.
Coherency scale.
To verify that the converted text retains the same content as the original text, we check the similarity in meaning between the text pairs using an LLM. Specifically, we create sample pairs from $\mathcal{B}$ and the converted subset $\mathcal{B}^{c}$, and inquire the LLM about the percentage of pairs that it considers to have the same meaning. This percentage is used as the coherency scale $s^{c}_{\text{coh}}$. Like the adversarialness scale, this value is rescaled to range between 0.1 and 1.
Finally, the selected style-conversion prompts are applied to the input texts of the entire dataset $\mathcal{D}$ to create augmented samples. These augmented samples are then used alongside the original samples to train the detector $f$. The detector is trained using binary cross-entropy loss. These prompt generation and selection processes are repeated over multiple training rounds.
4 Experiment
Table 1: Performance comparison with AdStyle in two different scenarios—Attack, where style-conversion attacks are performed, and Clean, where no attack is performed—across three fake news datasets. For the attack scenario, we report the average AUC of four style-conversion attacks. The best results are marked in bold. Our model demonstrates a significant performance improvement over all text-based fake news detectors in both style-conversion attack and clean scenarios.
Attack Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.6114 0.6904 0.8548 0.6920 0.7876 0.8453 0.8185 0.8674 0.8741
UDA 0.6241 0.7696 0.8564 0.7381 0.7865 0.8591 0.8615 0.9028 0.9297
RADAR 0.7218 0.7399 0.8571 0.7583 0.8170 0.8616 0.8535 0.8047 0.9086
ENDEF 0.6376 0.7579 0.8134 0.7405 0.7870 0.8615 0.8234 0.8950 0.8835
SheepDog 0.6525 0.8234 0.9009 0.7498 0.8357 0.8669 0.8926 0.9188 0.9630
AdStyle 0.7833 0.8919 0.9399 0.8134 0.8389 0.8721 0.9224 0.9531 0.9716
Clean Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.7393 0.8397 0.9355 0.7096 0.8104 0.8645 0.9311 0.9682 0.9892
UDA 0.7404 0.8783 0.9422 0.7422 0.8022 0.8666 0.9365 0.9724 0.9899
RADAR 0.7607 0.8495 0.9314 0.7593 0.8170 0.8630 0.9446 0.9773 0.9817
ENDEF 0.7776 0.8823 0.9294 0.7592 0.7991 0.8738 0.9234 0.9556 0.9871
SheepDog 0.7248 0.8229 0.9394 0.7490 0.8411 0.8641 0.9144 0.9459 0.9785
AdStyle 0.8996 0.9280 0.9460 0.8251 0.8493 0.8797 0.9509 0.9849 0.9889
Table 2: Performance comparison of ablations on the Politifact dataset. The results show the impact of style-conversion attacks using four different publishers (i.e., CNN, The New York Times, The Sun, and National Enquirer) and a clean scenario (i.e., Clean) where no attack is performed. Any modification or removal of model components leads to decreased performance.
Model CNN The New York Times The Sun National Enquirer Clean
Vanilla 0.8127 0.8687 0.8789 0.8591 0.9355
Random Selection 0.9075 0.9409 0.9355 0.9365 0.9412
Adv. only Selection 0.9035 0.9343 0.9318 0.9288 0.9405
w/o Uncertainty 0.9039 0.9333 0.9320 0.9402 0.9473
w/o Similarity 0.9193 0.9315 0.9297 0.9430 0.9409
w/o Score trajectory 0.8199 0.8917 0.9124 0.9066 0.9372
Full Components 0.9174 0.9444 0.9520 0.9460 0.9460
We evaluate the robustness of AdStyle under diverse style-conversion attacks across multiple datasets, comparing it with contemporary baselines. Then, we analyze the impact of model components on overall performance. Lastly, we conducted a qualitative analysis to investigate the characteristics of the style-conversion prompts generated by the LLM. Due to limited space, additional analysis, such as the performance evaluation over wider variety of paraphrasing attacks and comparisons with LLM-based zero-shot and in-context learning baselines, has been included in the Appendix.
4.1 Performance Evaluation
Dataset.
Our experiments use three real-world fake news benchmark datasets. We utilize PolitiFact and Gossipcop, drawn from the FakeNewsNet benchmark Shu et al. (2020), which focus on political claims and celebrity rumors, respectively. Additionally, we incorporate Constraint Felber (2021), a dataset specifically addressing COVID-19 related social media posts. Each dataset is randomly split into an 80% training set and a 20% test set. Detailed dataset statistics are provided in Table 3 of the Appendix.
Attack settings.
To assess robustness against style conversion attacks, we employ LLM-empowered techniques to reframe the test set using a variety of style conversion prompts, as illustrated in Figure 3. Following the original literature Wu and Hooi (2023), we use four well-known daily news sources as [publisher name]: CNN, The New York Times, The Sun, and National Enquirer. CNN and The New York Times are recognized for their reputable journalism, while The Sun and National Enquirer are characterized by their tabloid style. We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences, with the temperature set to 0 and the top-p value set to 1 by default. We conduct experiments using 10%, 25%, and 100% of the complete dataset to observe the effect of augmentation across different dataset sizes for training. AUC is used as our evaluation metric.
Rewrite the following article using the style of [publisher name]: [news article]
Figure 3: Prompt for style conversion. The “publisher name" part will be filled with the name of a representative publisher (e.g., newspaper or journal), and the “news article" part will contain the original news text.
Baselines.
We have implemented several existing text-based fake news detection strategies as baselines: (1) Vanilla, a text-based fake news detector using conventional binary cross entropy objective; (2) UDA, introduces consistency regularization objective between original text and diverse augmented variations. Xie et al. (2020); (3) RADAR, utilizes an adversarially trained paraphraser to generate augmented version of input sentences Hu et al. (2023); (4) ENDEF, mitigates entity bias in fake news data through causal learning Zhu et al. (2022); (5) SheepDog, introduces predefined style-conversion prompts to augment the styles of input text via LLM Wu and Hooi (2023). We follow the original paper’s setting and details for baseline implementations.
Implementation details.
All models are evaluated under uniform experimental conditions to ensure fair comparison. This consistency extends to the choice of backbone network, optimizer, and learning rate. We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences and measuring coherency, with the temperature set to 0 and the top-p value set to 1 by default. Our training process for the detector is conducted over 10 rounds, with one training epoch per round. When evaluating the style-conversion prompts, we randomly selected 30 samples from the training dataset to apply augmentation (i.e., $M=30$). In each round, 30 prompt candidates were generated by the LLM (i.e., $S=30$), from which 3 were chosen for augmentation (i.e., $k=3$) by our selection strategy. The training utilizes the AdamW optimizer with a learning rate of 1e-5 and a batch size of 8. For measuring diversity, the text embeddings are generated using a pre-trained BERT-based uncased model from HuggingFace Transformers. Two V100 GPUs were utilized for all experiments.
Result.
Table 1 compares the performance of detector algorithms in both clean scenarios, where no attack is performed, and adversarial scenarios, where style-conversion attacks are applied. Due to space limitations, the average AUC results under the four different style-conversion attacks are reported. Detailed results for each prompt can be found in the Appendix (See Table 6 in the Appendix). According to the results, AdStyle consistently outperforms the baselines across all cases, including both clean and adversarial scenarios. This demonstrates that our augmentation strategy enhances both the robustness and generalizability of the detector. Notably, our model’s effectiveness is more pronounced when compared to other baselines, especially with smaller datasets. Figure 4 shows the AUC for each style-conversion attack over different rounds. The performance gradually improves as the rounds progress, indicating that continually discovering adversarial augmentations is beneficial.
Figure 4: Performance changes across rounds on the PolitiFact dataset for four style-conversion attacks. The x-axis represents the training rounds, and the y-axis represents the detector’s AUC. For all attacks, the detector’s performance improved as the rounds progressed.
4.2 In-Depth Performance Analysis
We have demonstrated AdStyle’s effectiveness. In this section, we further examine the contribution of each component and provide a qualitative analysis on our adversarial style-conversion prompts.
Ablation study.
The proposed method integrates two main modules: adversarial style-conversion prompts generation and selection. To evaluate their individual contributions, we conduct experiments where we either remove each component or substituted it with an alternative within the full model. This results in six distinct configurations for analysis: (1) Full Components: Our complete method with all components; (2) Random Selection: The method that randomly select prompts from candidates instead of using our selection strategy; (3) Adversarial only Selection: The method that selects top-$k$ adversarial prompts, not considering diversity and coherence; (4) w/o Adversarialness: The method that omits the adversarialness score in our selection strategy. (5) w/o Coherency: The method that omits the similarity score in our selection strategy. (6) w/o Score trajectory: The method without score trajectory component for the style-conversion prompt generation.
Table 2 demonstrates that omitting any component leads to a decrease in performance for certain style conversions attack. While omitted due to space limitation, it can be confirmed that our sampling strategy converges faster than other sampling strategies (See Figure 8 in Appendix).
(a) Diversity
(b) Adversarialness
(c) Coherency
Figure 5: Qualitative analyses of AdStyle. (a) Histogram of diversity in embedding ($\mathbf{z}_{c}$) of conversion prompts sampled by our method and adversarial only selection. (b) Histogram of adversarialness scores ($s_{adv}^{c}$) for selected and unselected prompts (c) Histogram of coherency scores for selected and unselected prompts
Qualitative analyses.
First, we verified that our sampling strategy effectively selects diverse style-conversion prompts with high adversarialness and coherency. Figure 5(a) visualizes the diversity of prompts selected by our method compared to those chosen solely based on adversarialness scores (i.e., Adversarial-only Selection, the third model in our ablation study). We measure diversity using the average cosine similarity of every pair of selected prompts’ embeddings, $\mathbf{z}_{c}$ (Eq. 2):
$\text{Diveristy}(\mathcal{C})$ $=1-{1\over|\mathcal{C}|}\sum_{(c_{i},c_{j})\in\mathcal{C}}{\text{sim}(\mathbf{z}_{c_{i}},\mathbf{z}_{c_{j}})},$ (3)
where $\mathcal{C}$ is the set of prompt pairs, $\text{sim}(\cdot)$ represents the cosine similarity. The result in the Figure 5(a) indicate that our strategy results in a more diverse set of augmentations compared to selection based on adversarialness alone.
Figure 5(b) and 5(c) illustrate the Adversarialness and Coherency, respectively, of our selected prompts compared to remaining unselected prompts. Adversarialness was measured using the $s_{adv}^{c}$ score (Section 3.3), and coherency was calculated as the cosine similarity between the original text and its augmented version using semantic BERT embeddings Chanchani and Huang (2023). We can also observe that, for both metrics, prompts selected through AdStyle exhibit higher values compared to unselected prompts. This suggests that our strategy effectively selects prompts by considering both adversarialness and coherency.
Next, we qualitatively analyzed the characteristics of our style-conversion prompts. Figure 6 shows examples of prompts generated and selected by AdStyle. Previous work relied on prompts that were more conventional, such as “neutral" or “sensational." However, our prompts included creative phrases beyond typical human suggestions, which proved to be more adversarial to the detector. Currently, we use a fixed format for the initial set of prompts, which results in the generation of prompts with a similar format. We expect that using a more diverse initial set of prompts will allow the LLM to optimize prompts within a wider search space.
Round 0 P1: Rewrite the following article in a nonsensical and absurdly exaggerated tone with a hint of horror P2: Rewrite the following article in a sarcastic and mocking tone P3: Rewrite the following article in a chaotic and disorganized tone Round 1 P1: Rewrite the following article in a haunting and macabre tone with a sense of impending horror and madness P2: Rewrite the following article in a cryptic and enigmatic tone P3: Rewrite the following article in a malevolent and apocalyptic tone with a sense of impending doom and destruction, while also incorporating elements of surrealism and existential dread
Figure 6: Examples of our adversarial style-conversion prompts selected for the PolitiFact dataset
5 Conclusion
This paper presents a robust fake news detection method that withstands various style-conversion and paraphrasing attacks through adversarial style-conversion. Unlike traditional detectors that use predefined, agnostic augmentations, AdStyle employs tailored augmentations that shift samples in the direction of the detector’s current decision boundary using style-conversion prompts, functioning similarly to adversarial noise. Among the various prompt candidates generated by the LLM, we selected an efficient set of prompts for training by considering diversity, coherency, and adversarialness. As a result, we were able to train a detector that exhibits high robustness and generalizability against a wide range of attacks. We believe that our work helps filter fake news and contribute to a better exchange of information in online.
Appendix A APPENDIX
A.1 Data & Implementation Details
Dataset overview.
This paper utilizes a diverse range of datasets sourced from publicly available fake news detection resources, selected for their unique topical focuses. We utilize PolitiFact and Gossipcop, drawn from the FakeNewsNet benchmark Shu et al. (2020), addressing political claims and celebrity rumors respectively. Furthermore, we include Constraint Felber (2021), a dataset focused on COVID-19 related social media posts. Each dataset is randomly divided into an 80% training set and a 20% test set. Detailed statistics for these datasets can be found in Table 3.
Table 3: Statistics of fake news datasets.
Dataset PolitiFact GossipCop Constraint
# of News Articles 774 7,916 8,418
# of Real News 399 3,958 4,406
# of Fake News 375 3,958 4,012
Implementation details.
For consistency, experiments replicating existing baselines maintained fixed settings for learning, backbone network, and other relevant parameters. Augmented variations for UDA were generated using back-translation through German. RADAR employed adversarial training to learn paraphrases, utilizing the maximization of the binary cross-entropy loss of a fake news detector as a reward signal, with the T5-large model serving as the paraphraser. Named entities for ENDEF were extracted using the bert-base-NER model from Hugging Face. For SheepDog, augmentations were generated following the original paper’s methodology, using the prompt format illustrated in Figure 7 and four tones: "objective and professional," "neutral," "emotionally triggering," and "sensational". We also utilized the same four tones and prompt format as our predefined initial prompt set. Two V100 GPUs were utilized for all experiments.
Rewrite the following article in a/an [tone]: [news article]
Figure 7: Prompt for style-conversion. The “tone" part will be filled with the desired tone, and the “news article" part will contain the original news text.
A.2 Further Results on Evaluation.
Result on selection strategies across rounds.
Figure 8 shows the AUC for each selection strategy over different training rounds. Random Selection leads to slow convergence and unstable training, while Adversarial-only Selection, limited by its narrow focus on adversarial prompts, can negatively impact performance in later stages of training due to a lack of diversity in the generated prompts. Notably, without score trajectories, the model struggles to generate prompts that contribute effectively to the training process.
Figure 8: Performance of selection strategies over training rounds on the PolitiFact dataset. The x-axis represents the training rounds, while the y-axis represents the detector’s AUC.
Table 4: Performance comparison of different attack scenarios on the PolitiFact Dataset. (A: Adversarial prompt, B: Summarization prompt, C: In-Context prompt D: Adversarial Paraphraser)
Model A B C D
Vanilla 0.8881 0.8522 0.8896 0.8305
UDA 0.8624 0.8363 0.8924 0.8682
RADAR 0.9096 0.8754 0.9234 0.9007
ENDEF 0.8628 0.7978 0.9009 0.8682
SheepDog 0.9297 0.9205 0.9276 0.8995
Ours 0.9456 0.9416 0.9425 0.9212
Comparison on other possible attack scenarios.
We have conducted additional comparison experiments with more diverse attack scenarios: (1) Adversarial prompt: Given a news article and its label, the prompt instructs the LLM to rewrite the article to evade detection as the given label. (2) Summarization prompt: A prompt instructing the LLM to summarize the news article without incorporating stylistic elements. (3) In-Context prompt: A prompt providing an example of a recent, real CNN article, instructing the LLM to rewrite the given article in the same style. (4) Adversarial Paraphraser: The attack utilizes a paraphraser adversarially trained on the training dataset. The formats for each attack prompt are described in Figures 9 to 11. Based on the experimental results in Table 4, we confirm that our proposed model demonstrates a non-trivial performance improvement compared to other baselines against these attacks.
<EXAMPLES>: [news article] Determine whether the given <EXAMPLES> content is fact or fake, mentioning the style of the article that led to that decision following format: Answer: {Fact or Fake} Reason: {one-sentence reason }
Your current article is: [news article] But this article is detected as [Answer] due to the following [Reason]. Based on the above information, rewrite a new improved article not to be detected as [news label], maintaining the original content, as follows:
Figure 9: The format of Adversarial prompt. First, use the prompt above to extract the answer and reason of factuality of the given news article. Then, use the following prompt for paraphrasing to avoid detection based on the identified reason. The "news article" section contains the original news text, and the "news label" section contains the corresponding news label.
Summarize the following article, ensuring the content remains the same: [news article]
Figure 10: The format for Summarization prompt. The “news article" part will be filled with the original news text.
<EXAMPLES>: [news article example] Rewrite the following article as the writing style of <EXAMPLES> : [news article]
Figure 11: The format for In-Context prompt. The "news article example" section is filled with an example from a specific publisher, while the "news article" section contains the original news text.
Comparison with LLM-based Baselines.
We compared AdStyle with other LLM-based baselines: GPT-3.5 OpenAI (2022), GPT-4 OpenAI (2023), and In-Context Learning Brown et al. (2020). OpenAI APIs are utilized to predict news article labels, with the temperature set to 0 to ensure stable and consistent predictions. In the case of In-context Learning, one example each of fake news and real news was provided. The prompt formats for the LLM-based baselines are as follows:
Does the following contain real or fake news? Answer in one word with either ‘Real’or ‘Fake’: [news article]
Figure 12: Prompt for LLM-based baselines. The "news article" section contains the original news text.
Table 5: Performance comparison of different LLM-based baselines on the PolitiFact Dataset. (NY: The New York Times, TS: The Sun, NE: National Enquirer)
Model CNN NY TS NE Clean
GPT-3.5 0.5820 0.6242 0.6173 0.5274 0.7037
GPT-4 0.7133 0.6748 0.7025 0.6021 0.7823
In-Context 0.6954 0.6504 0.6875 0.5754 0.7383
Ours 0.9174 0.9444 0.9520 0.9460 0.9460
Full result on style-conversion attack.
Table 6 reports the AUC for each style-conversion attack. Paired sample t-tests were conducted to compare our method with the second-best baseline, Sheepdog, across different ratios for each dataset. Except for the 25% ratio in the Gossipcop dataset, all p-values were significantly smaller than 0.05, indicating statistically significant differences.
Table 6: Performance comparison with AdStyle in four different style—conversion attacks. For the attack scenario, we report the average AUC of four style-conversion attacks. The best results are marked in bold. Our model demonstrates a significant performance improvement over all text-based fake news detectors in diverse style-conversion attack scenarios. (NY: The New York Times, TS: The Sun, NE: National Enquirer )
CNN Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.5664 0.6566 0.8127 0.6841 0.7913 0.8518 0.8563 0.8175 0.8056
UDA 0.5906 0.7174 0.8009 0.7294 0.7859 0.8599 0.8305 0.8594 0.9021
RADAR 0.7205 0.7301 0.7926 0.7683 0.8138 0.8666 0.8099 0.8400 0.8873
ENDEF 0.5426 0.6679 0.7588 0.7288 0.7839 0.8624 0.7695 0.8666 0.8353
SheepDog 0.6305 0.8269 0.8683 0.7472 0.8393 0.8706 0.9200 0.9022 0.9544
AdStyle 0.7547 0.8822 0.9174 0.8195 0.8429 0.8744 0.9344 0.9430 0.9646
NY Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.5899 0.7334 0.8687 0.6750 0.7797 0.8465 0.8522 0.8939 0.8868
UDA 0.6189 0.7644 0.8744 0.7207 0.7795 0.8558 0.8745 0.9184 0.9297
RADAR 0.6966 0.7853 0.8586 0.7481 0.7960 0.8574 0.8701 0.8974 0.9176
ENDEF 0.6152 0.7630 0.8150 0.7227 0.7837 0.8609 0.8341 0.9067 0.8918
SheepDog 0.6310 0.7990 0.9090 0.7420 0.8309 0.8650 0.8937 0.9254 0.9645
AdStyle 0.7599 0.8940 0.9444 0.8118 0.8389 0.8686 0.9243 0.9594 0.9720
TS Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.6913 0.6859 0.8789 0.7068 0.7918 0.8430 0.7822 0.8727 0.8974
UDA 0.6646 0.8072 0.8883 0.7488 0.7904 0.8613 0.8662 0.9145 0.9446
RADAR 0.7271 0.7173 0.8909 0.7594 0.8062 0.8620 0.8564 0.8882 0.9132
ENDEF 0.7001 0.8132 0.8444 0.7603 0.7909 0.8639 0.8445 0.8955 0.9043
SheepDog 0.7070 0.8408 0.9096 0.7580 0.8395 0.8688 0.8962 0.9217 0.9665
AdStyle 0.8444 0.9002 0.9520 0.8117 0.8347 0.8751 0.9219 0.9521 0.9727
NE Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.6432 0.6856 0.8591 0.7020 0.7875 0.8399 0.7833 0.8853 0.9067
UDA 0.6557 0.7892 0.8622 0.7450 0.7902 0.8595 0.8747 0.9191 0.9423
RADAR 0.7428 0.7269 0.8864 0.7573 0.8026 0.8556 0.8775 0.8919 0.9161
ENDEF 0.6924 0.7877 0.8355 0.7502 0.7897 0.8589 0.8456 0.9112 0.9025
SheepDog 0.6636 0.8268 0.9166 0.7521 0.8331 0.8632 0.8606 0.9257 0.9667
AdStyle 0.8029 0.8913 0.9460 0.8166 0.8391 0.8702 0.9092 0.9579 0.9771

### v3

@@
 Adversarial Style Augmentation via Large Language Model for Robust Fake News Detection
-Abstract
-The spread of fake news negatively impacts individuals and is regarded as a significant social challenge that needs to be addressed.
-A number of algorithmic and insightful features have been identified for detecting fake news.
-However, with the recent LLMs and their advanced generation capabilities, many of the detectable features (e.g., style-conversion attacks) can be altered, making it more challenging to distinguish from real news.
-This study proposes adversarial style augmentation, AdStyle, to train a fake news detector that remains robust against various style-conversion attacks.
-Our model’s key mechanism is the careful use of LLMs to automatically generate a diverse yet coherent range of style-conversion attack prompts.
-This improves the generation of prompts that are particularly difficult for the detector to handle.
-Experiments show that our augmentation strategy improves robustness and detection performance when tested on fake news benchmark datasets.
-Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection
-Sungwon Park1, Sungwon Han1, Meeyoung Cha2 1School of Computing, KAIST, Daejeon, Republic of Korea 2MPI for Security and Privacy, Bochum, Germany {psw0416, lion4151}@kaist.ac.kr, mia.cha@mpi-sp.org
-1 Introduction
-With the widespread use of the internet and the emergence of various social media platforms, people have begun to freely share information they know and their own creative stories.
-However, the ease of sharing and consuming information beyond traditional news organizations has also significantly contributed to the spread of fake news Cha et al.
-(2021); Fisher et al.
-(2016); Treen et al.
-(2020).
-Fake news is created to provide readers with false information as propaganda or to guide public perception in a desired direction for intentional objectives Lazer et al.
-(2018); Yi et al.
-(2024).
-The spread of such false information has had profound negative impacts on individuals and society, becoming a major social challenge that needs to be addressed.
-Manually determining the authenticity of all information on the internet is extremely costly in terms of time and resources.
-Thus, prior advances have mainly focused on developing automated fake news detectors using machine learning techniques Pérez-Rosas et al.
-(2018); Reis et al.
-(2019); Shu et al.
-(2019).
-For example, several studies have extracted sentiment-related or political features commonly found in fake news, either through human-crafted Potthast et al.
-(2018); Rashkin et al.
-(2017) or data-driven methods Reis et al.
-(2019); Shu et al.
-(2019), and used them for detection.
-With the advent of language models (e.g., GPT, BERT) Kenton and Toutanova (2019); Radford et al.
-(2019), efforts have been made to learn the textual style of fake news by extracting sentence embeddings to train detectors Chen et al.
-(2021); Yang et al.
-(2022).
-While these extracted textual styles or human-crafted features have proven effective in identifying fake news, they allow attackers to bypass detection.
-In particular, the advent of large language models (LLMs) Chowdhery et al.
-(2022); OpenAI (2023) has made it possible to easily and automatically paraphrase sentences in a user-desired direction through prompts (e.g., "change the given text to an objective and professional style"), making it more difficult to distinguish between AI-generated fake news and real news (i.e., style-conversion attacks) Koenders et al.
-(2021); Zhou et al.
-(2019).
-Recent literature has proposed strategies to augment the styles of input text via some manually defined style-conversion prompts, which are taken by LLM for paraphrasing, but it remains challenging to address all types of prompts that attackers might use to deceive detectors Wu and Hooi (2023).
-In this study, we propose adversarial style augmentation, AdStyle, to train a robust fake news detector that can withstand various style-conversion attacks by attackers.
-In contrast to earlier work, which used predefined style-conversion prompts as detector-agnostic augmentations, AdStyle tries to find prompts for adversarial style augmentations that are specific to the detector, which makes predictions uncertain.
-This means that our augmentations add noise to the style features in the direction of the detector’s decision boundary, while maintaining the text’s content integrity.
-Specifically, to search for and optimize prompts for LLM that are adversarial to the detector, we introduce an automated prompt engineering technique Yang et al.
-(2023).
-By providing LLMs with style-conversion prompts and the detector’s performance under these augmentations, LLM can infer patterns between the prompts and performance, enabling the search for the most adversarial prompts.
-AdStyle proceeds as follows: To generate augmented samples, we first select a random subset of the dataset and apply a pool of style-conversion prompts to it.
-Then, each augmentation set created by different prompts is fed into the detector, and the score based on the AUC between the predictions and the ground-truth labels is measured.
-The prompt-score pairs are provided to the LLM, which uses this information to generate new style-conversion prompt candidates.
-From these candidates, we select the top-$k$ prompts that are diverse and make the detector’s predictions most uncertain without significantly altering the original content.
-These selected prompts are applied to the entire dataset and used for detector training.
-The chosen prompts are added to the prompt pool, which is repeated over several training rounds.
-We experimented with existing fake news benchmark datasets under various style-conversion attack scenarios.
-As a result, our augmentation strategy demonstrated higher robustness and detection performance compared to previous methodologies.
-Additionally, our approach is scalable to different attack strategies by adding new attack prompts to the style-conversion prompt pool.
-We plan to release our code after publication.
-2 Related Work
-2.1 Automated Detection of Fake News
-Manually detecting fake news among the vast amount of content on the internet is intractable.
-Consequently, many studies have focused on leveraging machine learning techniques to automatically detect fake news Pérez-Rosas et al.
-(2018); Reis et al.
-(2019); Shu et al.
-(2019).
-For example, supervised learning methodologies extract textual features from fake news texts using benchmark datasets and ground-truth labels Reis et al.
-(2019).
-These textual features can include deep features based on artificial neural networks Reis et al.
-(2019); Shu et al.
-(2019) as well as manually defined features such as sentiment and political bias Potthast et al.
-(2018); Rashkin et al.
-(2017).
-Other approaches include domain adaptation methods to enhance detection generalizability across various domains and topics Mosallanezhad et al.
-(2022); Nan et al.
-(2022), and knowledge-based methods that rely on external data to distinguish false information Dun et al.
-(2021).
-With the advent of LLMs, new methods have emerged that utilize LLMs’ prior knowledge to identify fake news Hu et al.
-(2024), including research on detecting machine-generated fake news Chen and Shu (2023).
-In this work, we aim to prevent malicious attackers from deceiving automated fake news detectors through text modifications such as paraphrasing.
-The proposed method is an augmentation strategy to enhance robustness against text style perturbations as an add-on to existing detection models.
-Our contribution is agnostic to the design of the detection model.
-Figure 1: Illustration of how AdStyle works.
-Our model consists of multiple rounds of training.
-In each round, the model uses the style-conversion prompt and prediction confusion score from the previous round to generate prompt candidates that can maximize the confusion of the detector model.
-Subsequently, considering aspects such as adversarialness, coherency, and diversity, a subset of the training dataset is used to select the top-$k$ prompts, which are finally used as augmentations in training.
-2.2 Attack on Fake News Detection
-To test the robustness of fake news detection, various attack methods have been studied Koenders et al.
-(2021); Zhou et al.
-(2019).
-These include injecting misinformation by changing the order of subjects and objects or causes and effects while maintaining the textual features used for detection Koenders et al.
-(2021).
-Other methods involve creating fact distortions by altering or exaggerating words related to people, time, or places while preserving the sentence structure Zhou et al.
-(2019).
-Additionally, approaches that use the text generation capabilities of LLMs to change the style of sentences have been proposed Wu and Hooi (2023).
-Our method aims to perform adversarial training of the detection model by generating augmentations that perturb the text’s style while preserving the content as much as possible.
-This approach reduces the impact of spurious style features on the detection model, making it more robust against such attacks.
-2.3 Prompt Engineering for LLM
-Training with extensive text corpora, LLMs have demonstrated their utility across various domain tasks Chowdhery et al.
-(2022); OpenAI (2023); Pang et al.
-(2024).
-To better harness the prior knowledge and reasoning abilities of LLMs, strategies for providing appropriate input prompts have also been researched.
-For instance, in-context learning methods involve providing examples of the desired input-output format to guide the LLM in producing the correct output Dong et al.
-(2022).
-Additionally, engineering techniques such as chain-of-thought prompting, which includes additional reasoning steps in the responses, have been developed to enhance the reasoning abilities of LLMs Wei et al.
-(2022).
-Recently, methods for automatically finding the most suitable prompts for a given task, known as automated prompt engineering, have shown promising results Yang et al.
-(2023); Zhou et al.
-(2022).
-In this paper, we adopt automated prompt engineering techniques to identify adversarial prompts that most effectively confuse the fake news detector.
-This process replaces the traditional gradient descent method used in the image domain to find adversarial noise, enabling the creation of adversarially augmented versions of input text via LLM.
-3 Method
-3.1 Overview
+Abstract.
+The spread of fake news harms individuals and presents a critical social challenge that must be addressed.
+Although numerous algorithmic and insightful features have been developed to detect fake news, many of these features can be manipulated with style-conversion attacks, especially with the emergence of advanced language models, making it more difficult to differentiate from genuine news.
+This study proposes adversarial style augmentation, AdStyle, designed to train a fake news detector that remains robust against various style-conversion attacks.
+The primary mechanism involves the strategic use of LLMs to automatically generate a diverse and coherent array of style-conversion attack prompts, enhancing the generation of particularly challenging prompts for the detector.
+Experiments indicate that our augmentation strategy significantly improves robustness and detection performance when evaluated on fake news benchmark datasets.
+Misinformation, Adversarial Training, Large Language Model
+copyright: acmlicensedjournalyear: 2025conference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714569isbn: 979-8-4007-1274-6/25/04ccs: Security and privacy Software and application securityccs: Computing methodologies Artificial intelligence
+1. Introduction
+In today’s digital landscape, people seek information through various channels, including social media.
+News content continues to hold substantial importance on these platforms, significantly shaping individuals’ thoughts and decisions.
+However, the ease of sharing and consuming information outside traditional news outlets has introduced several challenges.
+Among them is the rise of alternative media and news-like content, which can contain completely or partially false information (Vosoughi et al., 2018).
+The spread of such misinformation has negative impacts on individuals and is considered a major social challenge that requires urgent attention (Cha et al., 2021).
+Given the vast amount of information shared on social platforms, manually detecting fake news has become impractical.
+Consequently, these efforts are increasingly supported and managed by algorithms.
+For example, sentiment and topic features, as well as the temporal and structural patterns of diffusion networks, have proven effective in detecting fake news (Kwon et al., 2013; Potthast et al., 2018; Rashkin et al., 2017).
+These patterns can be utilized by machine learning and deep learning (Ma et al., 2016; Shu et al., 2019; Pérez-Rosas et al., 2018).
+Large language models (LLMs) have also been used to learn the textual style of fake news by extracting sentence embeddings to train detectors (Augenstein et al., 2024b; Hu et al., 2024; Chen and Shu, 2023).
+Although certain textual styles or linguistic cues can indicate fake news, attackers can circumvent such detection by rearranging the order of subjects and objects or by imitating specific styles (Koenders et al., 2021; Zhou et al., 2019).
+LLMs can easily paraphrase sentences in any desired manner through prompts (e.g., ”change this text into a NYTimes style”), making it challenging to differentiate AI-generated news from authentic ones.
+These manipulations, referred to as style-conversion attacks, significantly complicate the task of maintaining information veracity in the digital media landscape (Koenders et al., 2021; Zhou et al., 2019).
+In this study, we present AdStyle, an adversarial style augmentation method designed to train a fake news detector that can withstand various style-conversion attacks.
+Unlike previous methods that relied on predefined style-conversion prompts, our approach identifies prompts to generate adversarial style augmentations tailored to a specific detector.
+This process adds noise to the style features within the detector’s decision boundary, while preserving content-wise integrity.
+To customize prompts that are adversarial to the detector, we implement an automated prompt engineering technique, similar to (Yang et al., 2023).
+By providing style-conversion prompts and evaluating the detector’s performance under these augmentations, the LLM can discern patterns between prompts and performance, facilitating search for the most effective prompts.
+When evaluated on fake news benchmarks such as PolitiFact, GossipCop, and Constraint under various style-conversion attack scenarios, our augmentation strategy exhibited enhanced robustness and detection performance compared to current state-of-the-art methods.
+Our strategy preserves the content of the sentence while altering its structure to increase perplexity according to the LLM-based detector.
+This adjustment results in a sentence structure that the LLM has not frequently encountered during its pre-training phase, effectively deterring malicious style-conversion attacks, such as paraphrasing.
+Additionally, our method serves as an augmentation strategy that can be integrated with any existing detection models, regardless of their design.
+We release the code and implementation details of our fake news detection against style-conversion attacks for the broader use and greater impact within the research community and industry: https://github.com/deu30303/AdStyle.
+2. Related Work
+2.1. Automated Detection of Fake News
+Among the machine-driven fake news detection methods are learning methodologies extract textual features from fake news texts using benchmark datasets and ground-truth labels (Reis et al., 2019).
+These textual features can include deep features based on artificial neural networks (Shu et al., 2019) as well as manually defined features such as sentiment and political bias (Kwon et al., 2013; Potthast et al., 2018; Rashkin et al., 2017).
+Other approaches include domain adaptation methods to enhance detection generalizability across various domains and topics (Mosallanezhad et al., 2022; Nan et al., 2022), and knowledge-based methods that rely on external data to distinguish false information (Dun et al., 2021).
+With the advent of LLMs, new methods have emerged that utilize LLMs’ prior knowledge to identify fake news (Hu et al., 2024), including research on detecting machine-generated fake news (Chen and Shu, 2023).
+One approach involves using LLMs to generate synthetic reactions and comments on news articles from diverse perspectives, which are then leveraged for fake news detection (Wan et al., 2024).
+Furthermore, LLM-generated logic has been employed to enhance the interpretability of fake news detection models (Liu et al., 2024).
+Despite these advancements, fake news detection using LLMs remains a challenging task due to factuality hallucinations, which can undermine reliability (Augenstein et al., 2024a).
+2.2. Attack on Fake News Detection
+Various attack methods have been studied to test the robustness of fake news detection (Koenders et al., 2021; Zhou et al., 2019).
+These methods include injecting misinformation by rearranging subjects and objects or causes and effects while retaining the textual features used for detection (Koenders et al., 2021).
+Other techniques involve creating fact distortions by modifying or exaggerating words related to people, time, or places, while preserving sentence structure (Zhou et al., 2019).
+Another line of research simulates the behavior of malicious users on social media by manipulating news article comments or employing a multi-agent reinforcement learning framework to generate adversarial content (Le et al., 2020; Wang et al., 2023).
+Recently, methods leveraging the text generation capabilities of LLMs to modify the writing style of sentences have also been proposed (Wu and Hooi, 2023).
+In this study, we propose an adversarial training method to mitigate style-conversion attacks in fake news detection.
+Our approach utilizes automated prompt engineering to identify the most effective style-conversion prompts for a given fake news dataset (Yang et al., 2023; Zhou et al., 2022).
+Figure 1.
+Illustration of AdStyle.
+The model undergoes multiple rounds of training.
+In each round, it utilizes the style-conversion prompt and prediction confusion score from the previous round to create prompt candidates aimed at maximizing the detector model’s confusion.
+Subsequently, a subset of the training dataset is used to select the top-$k$ prompts, based on factors such as adversarialness, coherency, and diversity.
+These selected prompts are then employed as augmentations in the training process.
+Repetition of this process allows the model to consider a wide range of adversarial prompts, making the director more resilient to future attacks.
+3. Method
+3.1. Overview
 Let $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$ be a dataset containing news $\mathbf{d}_{i}$ and the corresponding ground-truth binary veracity label $y_{i}$ (indicating whether the news is true or fake).
 Each news item $\mathbf{d}_{i}$ is composed of natural language-based text.
-This study aims to train a language model based fake news detector $f$ using the labeled dataset to predict the veracity labels.
-Our main goal is to develop a detector that remains robust even when an attacker perturbs the textual style, such as the order and format, while preserving the meaning of the sentences.
-Figure 1 illustrates how our model works.
-AdStyle generates style-conversion prompts and performs augmentation over multiple rounds.
-Each round includes a following process.
+Our goal is to train a fake news detector $f$ based on the labeled dataset and predict veracity labels.
+In doing so, we ensure that a detector remains robust even when an attacker perturbs the textual style, such as the order and format, while preserving the meaning of the sentences.
+Figure 1 illustrates the workflow.
+Our method, called AdStyle, generates style-conversion prompts and performs augmentation over multiple rounds of the following process.
 Firstly, leveraging the reasoning ability of LLM, adversarial style-conversion prompt candidates are generated.
 These style-conversion prompts contain instructions on how to transform the given sentences and are used as inputs to the LLM along with the original sentences for conversion.
-Inspired by automated prompt engineering Yang et al.
-(2023), the style-conversion prompts and detector prediction score pairs used in previous rounds are included as in-context demonstrations to guide the LLM in searching prompt candidates that maximally confuse the detector (Section 3.2).
+Following the automated prompt engineering (Yang et al., 2023), the style-conversion prompts and detector prediction score pairs used in previous rounds are included as in-context demonstrations to guide the LLM in searching prompt candidates that maximally confuse the detector (Section 3.2).
 Next, a subset of the dataset is selected, and these discovered candidates are applied to perform conversions.
@@
 We describe each step’s details below.
-3.2 Generating Adversarial Style-Conversion Prompts with LLM
-The style-conversion prompts we aim to generate are instructions that perturb only the textual style, such as the structure or format of the sentence, while preserving the content of the given input text.
-For example, an instruction like “Rewrite the following article in an objective and professional tone” can be used to transform the stylistic features of fake news to resemble real news.
-However, if such instructions are heuristically defined, they may not align with the actual decision boundary of the detector, reducing the training efficiency.
-Furthermore, if the instructions remain fixed throughout the training, the detector may memorize these conversion patterns, resulting in overfitting.
-Therefore, instead of predefining and fixing the instructions for conversion, we aim to find adversarial conversion prompts that add noise in the direction of the decision boundary of current state of detector, maximizing the confusion in the detector’s predictions.
-This approach is similar to adversarial training commonly used in the computer vision domain Madry et al.
-(2018), which enhances robustness against slight perturbations in the input.
+Now you will help me minimize a fake news detector performance value with style transfer attack instructions.
+I have a set of style transfer attack instructions that I will use to minimize the fake news detector’s performance.
+The pairs are arranged in descending order based on their function values, where lower values are better.
+Input: [Style-Conversion Prompt Example]
+Score: [Prediction Confusion Score]
+Input: … …
+Give me a new style transfer attack instruction that is different from all pairs above and has a lower performance value.
+Figure 2. Prompt for generating adversarial style-conversion prompts. The text in blue represents the score trajectory, and the remaining text represents the problem description.
+3.2. Adversarial Style-Conversion Prompts
+The style-conversion prompts we generate are instructions that perturb only the textual style, such as the structure or format of the sentence, while preserving the content of the given input text.
+For example, an instruction like “Rewrite the following article in an objective and professional tone” can transform the stylistic features of fake news to resemble authentic news.
+However, if these instructions are heuristically defined, they may not align with the detector’s actual decision boundary, thereby reducing training efficiency.
+Furthermore, if the instructions remain fixed throughout the training process, the detector may memorize these conversion patterns, leading to overfitting.
+Instead of predefining and fixing the instructions for conversion, we identify adversarial conversion prompts that introduce noise directed toward the decision boundary of the current detector, thereby maximizing prediction confusion.
+This approach is similar to adversarial training commonly used in the computer vision domain (Madry et al., 2018), which enhances robustness against slight input perturbations.
 However, for textual data, the discrete nature of the input makes it challenging to add noise directly through the detector’s gradient as in the image domain.
-To address this, we measure the confusion in prediction by conversion prompts (i.e., prediction confusion score), and feed this prompt-score pair to LLM for generating adversarial prompts, motivated from automated prompt engineering techniques.
-Now you will help me minimize a fake news detector performance value with style transfer attack instruction.
-I have some style transfer attack instructions and the fake news detector performance value using those instructions.
-The pairs are arranged in descending order based on their function values, where lower values are better.
-Input: [Style-Conversion Prompt Example] Score: [Prediction Confusion Score] Input: … … Give me a new style transfer attack instruction that is different from all pairs above, and has a performance value lower than any of the above.
-Figure 2: Prompt for generating adversarial style-conversion prompts. The text in blue represents the score trajectory component, while the remaining text represents the problem description component.
-Figure 2 shows an example of the LLM input used to generate adversarial style-conversion prompts. The LLM input consists of the problem description and the score trajectory components.
+To address this, we measure prediction confusion using conversion prompts (i.e., prediction confusion score) and feed this prompt-score pair to the LLM for generating adversarial prompts, inspired by automated prompt engineering techniques.
+Figure 2 shows an example of the LLM input used to generate adversarial style-conversion prompts.
+The LLM input consists of the problem description and the score trajectory components.
+We describe each component in turn.
 Problem description component.
@@
 Score trajectory component.
-Previous work has shown that LLMs can learn patterns from in-context demonstrations provided as input Dong et al.
-(2022); Yang et al.
-(2023).
-This component leverages this ability by providing previous round’s style-conversion prompts and their corresponding prediction confusion scores in the form of in-context demonstrations [^1] In the first round, a predefined set of prompts is used..
-The score is measured by selecting a subset $\mathcal{B}$ from the entire training dataset $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$, applying a conversion prompt $c$ to create a new set $\mathcal{B}^{c}=\{(\mathbf{d}^{c}_{i},y_{i})\}_{i=1}^{M}$, where $\mathbf{d}^{c}_{i}=\text{Convert}(c,\mathbf{d}_{i}),N>>M$, and then measuring the AUC score between the predictions and the ground-truth labels when $\mathcal{B}^{c}$ is fed into the detector.
+Previous research has demonstrated that LLMs can learn patterns from in-context demonstrations provided as input (Dong et al., 2022; Yang et al., 2023).
+The score trajectory component leverages this ability by providing style-conversion prompts from previous rounds and their corresponding prediction confusion scores in the form of in-context demonstrations.
+In the first round, a predefined set of prompts is excerpted from (Wu and Hooi, 2023).
+The score is calculated by selecting a subset $\mathcal{B}$ from the entire training dataset $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$, applying a conversion prompt $c$ to create a new set $\mathcal{B}^{c}=\{(\mathbf{d}^{c}_{i},y_{i})\}_{i=1}^{M}$, where $\mathbf{d}^{c}_{i}=\text{Convert}(c,\mathbf{d}_{i}),N>>M$, and then measuring the AUC score between the predictions and the ground-truth labels when $\mathcal{B}^{c}$ is fed into the detector.
 Specifically, the score of a conversion prompt $s_{c}$ is defined as:
-$s_{c}=|0.5-\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})|.$ (1)
+(Eq. 1) $s_{c}=|0.5-\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})|.$
 A lower score indicates a higher level of prediction confusion, implying that the conversion prompt has caused the detector’s predictions to become random with respect to the labels.
-Finding conversion prompts that cause high confusion (i.e., low score) suggests that the detector has not yet learned to handle those stylistic features, and the conversion prompts have placed the samples near the detector’s decision boundary, making them difficult to distinguish.
-By showing the LLM these in-context demonstrations, it can generate conversion prompts that differ from past ones while maximizing confusion (i.e., minimizing the score).
-Note that we avoid selecting conversion prompts that flip the detector’s original predictions (i.e., AUC <<much-less-than$<$ < 0.5), as these often disrupt the content along with the stylistic features, which is undesirable.
+Finding conversion prompts that lead to high confusion (i.e., low score) suggests that the detector has not yet learned to handle those stylistic features, and the conversion prompts have placed the samples near the detector’s decision boundary, making them difficult to distinguish.
+By providing these in-context demonstrations, the LLM can generate conversion prompts that differ from previous ones while maximizing confusion (i.e., minimizing the score).
+It is important to avoid selecting conversion prompts that flip the detector’s original predictions (i.e., AUC <<much-less-than$<$ < 0.5), as such cases will disrupt the content and stylistic features.
 We extract $S$ style-conversion prompts at a time using the input described above.
-3.3 Selecting Top-$k$ Adversarial Prompts
+3.3. Selecting Top-$k$ Adversarial Prompts
 Using all the style-conversion prompt candidates generated by the LLM can be computationally intensive, and not all prompts may be suitable for augmentation.
 For example, the set of style-conversion prompts used for augmentation should each provide adversarial perturbations that confuse the detector (i.e., adversarialness).
-Additionally, while altering the textual style, the prompts should not change the meaning of the sentences to prevent label noise (i.e., coherency).
-Furthermore, the more diverse the set of conversion directions covered by the prompts, the more efficient the augmentation process (i.e., diversity).
+While altering the textual style, the prompts should not change the meaning of the sentences to prevent label noise (i.e., coherency).
+The more diverse the set of conversion directions covered by the prompts, the more efficient the augmentation process (i.e., diversity).
 To select a set of style-conversion prompts that satisfies these three criteria, we propose a selection strategy.
@@
 We then compute the average embedding vectors for each subset and calculate the vector difference $\mathbf{z}_{c}$.
-$\mathbf{z}$ $=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{d}_{i}\in\mathcal{B}}{g(\mathbf{d}_{i})},\ \ \ \mathbf{z}\textquoteright=\frac{1}{|\mathcal{B}\textquoteright|}\sum_{\mathbf{d}^{c}_{i}\in\mathcal{B}\textquoteright}{g(\mathbf{d}^{c}_{i})}$
-$\mathbf{z}_{c}$ $=\mathbf{z}\textquoteright-\mathbf{z}$ (2)
+$\mathbf{z}$ $=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{d}_{i}\in\mathcal{B}}{g(\mathbf{d}_{i})},\ \ \ \mathbf{z}’=\frac{1}{|\mathcal{B}’|}\sum_{\mathbf{d}^{c}_{i}\in\mathcal{B}’}{g(\mathbf{d}^{c}_{i})}$
+(Eq. 2) $\mathbf{z}_{c}$ $=\mathbf{z}’-\mathbf{z}$
 Here, $\mathbf{z}_{c}$ represents the average change in embedding direction due to the conversion prompt $c$.
 We then calculate the adversarialness scale $s^{c}_{\text{adv}}$ and coherency scale $s^{c}_{\text{coh}}$ for each conversion prompt $c$, and rescale $\mathbf{z}_{c}$ accordingly (i.e., $\mathbf{\hat{z}}_{c}=\mathbf{z}_{c}\times s^{c}_{\text{adv}}\times s^{c}_{\text{coh}}$).
-Finally, we use the $k$-means++ initialization method Arthur et al.
-(2007) on these rescaled vectors to select $k$ prompts.
-The $k$-means++ initialization method helps select a diverse set of prompts that are adversarial and coherent, by choosing samples that are as far apart as possible Ash et al.
-(2019).
+Finally, we use the $k$-means++ initialization method (Arthur et al., 2007) on these rescaled vectors to select $k$ prompts.
+The $k$-means++ initialization method helps select a diverse set of prompts that are adversarial and coherent, by choosing samples that are as far apart as possible (Ash et al., 2019).
 The details of each scale are described below.
-Adversarialness scale.
-To measure how adversarial a given style-conversion prompt is to the detector, we define a new adversarialness scale similar to the confusion score defined in the previous section.
-Given the converted batch $\mathcal{B}^{c}$ by conversion prompt $c$, the adversarialness scale $s^{c}_{\text{adv}}$ is defined as $-1.8\cdot|\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})-0.5|+1$.
-This value increases as the AUC approaches 0.5, indicating that the prediction is more random.
-The coefficient 1.8 ensures that the scale ranges between 0.1 and 1, preventing it from being zero.
-Coherency scale.
+Adversarialness scale ($s^{c}_{\text{adv}}$).
+To measure how adversarial a given style-conversion prompt is to the detector, we newly define the adversarialness scale similar to the confusion score defined in the previous section.
+Given the converted batch $\mathcal{B}^{c}$ by conversion prompt $c$, the adversarialness scale $s^{c}_{\text{adv}}$ is defined as:
+(Eq. 3) $s^{c}_{\text{adv}}=-1.8\cdot|\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})-0.5|+1.$
+This value increases as the AUC approaches 0.5, indicating that the prediction is more random. The coefficient 1.8 ensures that the scale ranges between 0.1 and 1, preventing it from being zero.
+Coherency scale ($s^{c}_{\text{coh}}$).
 To verify that the converted text retains the same content as the original text, we check the similarity in meaning between the text pairs using an LLM.
-Specifically, we create sample pairs from $\mathcal{B}$ and the converted subset $\mathcal{B}^{c}$, and inquire the LLM about the percentage of pairs that it considers to have the same meaning.
+We create sample pairs from $\mathcal{B}$ and the converted subset $\mathcal{B}^{c}$, and inquire the LLM about the percentage of pairs that it considers to have the same meaning.
 This percentage is used as the coherency scale $s^{c}_{\text{coh}}$.
 Like the adversarialness scale, this value is rescaled to range between 0.1 and 1.
-Finally, the selected style-conversion prompts are applied to the input texts of the entire dataset $\mathcal{D}$ to create augmented samples.
+Finally, the selected style-conversion prompts via our score and $k$-means++ initialization method are applied to the input texts of the entire dataset $\mathcal{D}$ to create augmented samples.
 These augmented samples are then used alongside the original samples to train the detector $f$.
@@
 These prompt generation and selection processes are repeated over multiple training rounds.
-4 Experiment
-Table 1: Performance comparison with AdStyle in two different scenarios—Attack, where style-conversion attacks are performed, and Clean, where no attack is performed—across three fake news datasets.
+4. Experiment
+We evaluate the robustness of AdStyle under diverse style-conversion attacks across multiple datasets, comparing it with contemporary baselines.
+We then analyze the impact of various model components on overall performance.
+A comprehensive evaluation is also performed on a broader range of paraphrasing attacks, including comparisons with LLM-based zero-shot and in-context learning baselines.
+Lastly, a qualitative analysis is conducted to explore the characteristics of the generated style-conversion prompts.
+4.1. Performance Evaluation
+Dataset.
+We use three real-world fake news benchmark datasets: (1) PolitiFact and (2) Gossipcop, released by FakeNewsNet benchmark (Shu et al., 2020), which focus on political claims and celebrity rumors, respectively, and (3) Constraint (Felber, 2021), a dataset of social media posts on COVID-19.
+Datasets were split into an 80% training set and a 20% test set, as described in Table 1.
+Table 1. Statistics of fake news datasets.
+Dataset PolitiFact GossipCop Constraint
+# of News Articles 774 7,916 8,418
+# of Real News 399 3,958 4,406
+# of Fake News 375 3,958 4,012
+Table 2.
+Performance comparison with AdStyle in two different scenarios—Attack, where style-conversion attacks are performed, and Clean, where no attack is performed—across three fake news datasets.
 For the attack scenario, we report the average AUC of four style-conversion attacks.
 The best results are marked in bold.
+The values 0.1, 0.25, and 1 indicate the proportion of the dataset used for training relative to the full dataset.
 Our model demonstrates a significant performance improvement over all text-based fake news detectors in both style-conversion attack and clean scenarios.
@@
 AdStyle 0.8996 0.9280 0.9460 0.8251 0.8493 0.8797 0.9509 0.9849 0.9889
-Table 2: Performance comparison of ablations on the Politifact dataset.
+Attack settings.
+To assess robustness against style conversion attacks, we employ LLM-empowered techniques to reframe the test set using a variety of style conversion prompts, as illustrated in Figure 3.
+Following the original literature (Wu and Hooi, 2023), we use four prominent daily news sources as [publisher name]: CNN, The New York Times, The Sun, and National Enquirer.
+CNN and The New York Times were chosen as representative news outlets recognized for their reputable journalism, while The Sun and National Enquirer are characterized by their tabloid style.
+We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences, with the temperature set to 0 and the top-p value set to 1 by default.
+Experiments are conducted using 10%, 25%, and 100% of the complete dataset to observe the impact of augmentation on training performance, measured by AUC, across various dataset sizes.
+Rewrite the following article using the style of [publisher name]: [news article]
+Figure 3.
+Prompt for style conversion.
+The “publisher name” part will be filled with the name of a representative publisher (e.g., newspaper or journal), and the “news article” part will contain the original news text.
+Table 3.
+Performance comparison of ablations on the Politifact dataset.
 The results show the impact of style-conversion attacks using four different publishers (i.e., CNN, The New York Times, The Sun, and National Enquirer) and a clean scenario (i.e., Clean) where no attack is performed.
@@
 Random Selection 0.9075 0.9409 0.9355 0.9365 0.9412
-Adv. only Selection 0.9035 0.9343 0.9318 0.9288 0.9405
-w/o Uncertainty 0.9039 0.9333 0.9320 0.9402 0.9473
-w/o Similarity 0.9193 0.9315 0.9297 0.9430 0.9409
+Class Prompt 0.9131 0.9272 0.9333 0.9313 0.9471
+Adversarial only Selection 0.9035 0.9343 0.9318 0.9288 0.9405
+w/o Adversarialness 0.9039 0.9333 0.9320 0.9402 0.9473
+w/o Coherency 0.9193 0.9315 0.9297 0.9430 0.9409
 w/o Score trajectory 0.8199 0.8917 0.9124 0.9066 0.9372
 Full Components 0.9174 0.9444 0.9520 0.9460 0.9460
-We evaluate the robustness of AdStyle under diverse style-conversion attacks across multiple datasets, comparing it with contemporary baselines.
-Then, we analyze the impact of model components on overall performance.
-Lastly, we conducted a qualitative analysis to investigate the characteristics of the style-conversion prompts generated by the LLM.
-Due to limited space, additional analysis, such as the performance evaluation over wider variety of paraphrasing attacks and comparisons with LLM-based zero-shot and in-context learning baselines, has been included in the Appendix.
-4.1 Performance Evaluation
-Dataset.
-Our experiments use three real-world fake news benchmark datasets.
-We utilize PolitiFact and Gossipcop, drawn from the FakeNewsNet benchmark Shu et al.
-(2020), which focus on political claims and celebrity rumors, respectively.
-Additionally, we incorporate Constraint Felber (2021), a dataset specifically addressing COVID-19 related social media posts.
-Each dataset is randomly split into an 80% training set and a 20% test set.
-Detailed dataset statistics are provided in Table 3 of the Appendix.
-Attack settings.
-To assess robustness against style conversion attacks, we employ LLM-empowered techniques to reframe the test set using a variety of style conversion prompts, as illustrated in Figure 3.
-Following the original literature Wu and Hooi (2023), we use four well-known daily news sources as [publisher name]: CNN, The New York Times, The Sun, and National Enquirer.
-CNN and The New York Times are recognized for their reputable journalism, while The Sun and National Enquirer are characterized by their tabloid style.
-We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences, with the temperature set to 0 and the top-p value set to 1 by default.
-We conduct experiments using 10%, 25%, and 100% of the complete dataset to observe the effect of augmentation across different dataset sizes for training.
-AUC is used as our evaluation metric.
-Rewrite the following article using the style of [publisher name]: [news article]
-Figure 3: Prompt for style conversion.
-The “publisher name" part will be filled with the name of a representative publisher (e.g., newspaper or journal), and the “news article" part will contain the original news text.
 Baselines.
-We have implemented several existing text-based fake news detection strategies as baselines: (1) Vanilla, a text-based fake news detector using conventional binary cross entropy objective; (2) UDA, introduces consistency regularization objective between original text and diverse augmented variations.
-Xie et al.
-(2020); (3) RADAR, utilizes an adversarially trained paraphraser to generate augmented version of input sentences Hu et al.
-(2023); (4) ENDEF, mitigates entity bias in fake news data through causal learning Zhu et al.
-(2022); (5) SheepDog, introduces predefined style-conversion prompts to augment the styles of input text via LLM Wu and Hooi (2023).
+We implemented several existing text-based fake news detection strategies as baselines: (1) Vanilla, a text-based fake news detector using conventional binary cross entropy objective; (2) UDA, introduces consistency regularization objective between original text and diverse augmented variations.
+(Xie et al., 2020); (3) RADAR, utilizes an adversarially trained paraphraser to generate augmented version of input sentences (Hu et al., 2023); (4) ENDEF, mitigates entity bias in fake news data through causal learning (Zhu et al., 2022); (5) SheepDog, introduces predefined style-conversion prompts to augment the styles of input text via LLM (Wu and Hooi, 2023).
 We follow the original paper’s setting and details for baseline implementations.
@@
 We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences and measuring coherency, with the temperature set to 0 and the top-p value set to 1 by default.
-Our training process for the detector is conducted over 10 rounds, with one training epoch per round.
+The training process for the detector is conducted over 10 rounds, with one training epoch per round.
 When evaluating the style-conversion prompts, we randomly selected 30 samples from the training dataset to apply augmentation (i.e., $M=30$).
@@
 Two V100 GPUs were utilized for all experiments.
+Figure 4.
+Performance changes across rounds on the PolitiFact dataset for four different style-conversion attacks.
+The x-axis represents the training rounds, and the y-axis represents the detector’s AUC.
+For all attacks, the detector’s performance improved as the rounds progressed.
 Result.
-Table 1 compares the performance of detector algorithms in both clean scenarios, where no attack is performed, and adversarial scenarios, where style-conversion attacks are applied.
+Table 2 shows results comparing the performance of detector algorithms in both clean scenarios, where no attack is performed, and adversarial scenarios, where style-conversion attacks are applied.
 Due to space limitations, the average AUC results under the four different style-conversion attacks are reported.
-Detailed results for each prompt can be found in the Appendix (See Table 6 in the Appendix).
-According to the results, AdStyle consistently outperforms the baselines across all cases, including both clean and adversarial scenarios.
+We find that AdStyle consistently outperforms the baselines, including both clean and adversarial scenarios.
 This demonstrates that our augmentation strategy enhances both the robustness and generalizability of the detector.
-Notably, our model’s effectiveness is more pronounced when compared to other baselines, especially with smaller datasets.
+The model is particularly effective when compared to other baselines and for smaller datasets.
 Figure 4 shows the AUC for each style-conversion attack over different rounds.
 The performance gradually improves as the rounds progress, indicating that continually discovering adversarial augmentations is beneficial.
-Figure 4: Performance changes across rounds on the PolitiFact dataset for four style-conversion attacks.
-The x-axis represents the training rounds, and the y-axis represents the detector’s AUC.
-For all attacks, the detector’s performance improved as the rounds progressed.
-4.2 In-Depth Performance Analysis
-We have demonstrated AdStyle’s effectiveness. In this section, we further examine the contribution of each component and provide a qualitative analysis on our adversarial style-conversion prompts.
-Ablation study.
+4.2. Component Analysis
+We now examine the contribution of each component on our adversarial style-conversion prompts.
 The proposed method integrates two main modules: adversarial style-conversion prompts generation and selection.
 To evaluate their individual contributions, we conduct experiments where we either remove each component or substituted it with an alternative within the full model.
-This results in six distinct configurations for analysis: (1) Full Components: Our complete method with all components; (2) Random Selection: The method that randomly select prompts from candidates instead of using our selection strategy; (3) Adversarial only Selection: The method that selects top-$k$ adversarial prompts, not considering diversity and coherence; (4) w/o Adversarialness: The method that omits the adversarialness score in our selection strategy.
-(5) w/o Coherency: The method that omits the similarity score in our selection strategy.
-(6) w/o Score trajectory: The method without score trajectory component for the style-conversion prompt generation.
+This results in six distinct configurations for analysis: (1) Full Components: Our complete method with all components; (2) Random Selection: The method that randomly select conversion prompts from candidates instead of using our selection strategy (Sec.
+3.3); (3) Class Prompt: The method categorizes the confusion scores of conversion prompts into three levels: high, medium, or low.
+These categorized labels are then used as in-context demonstrations for generating adversarial style-conversion prompts (Sec.
+3.2), instead of relying on continuous confusion scores $s_{c}$ (Eq.
+1); (4) Adversarial only Selection: The method that selects top-$k$ adversarial prompts, not considering diversity and coherence criteria; (5) w/o Adversarialness: The method that omits the adversarialness criterion in our selection strategy; (6) w/o Coherency: The method that omits the coherency criterion in our selection strategy; (7) w/o Score trajectory: The method without score trajectory component for the style-conversion prompt generation (Sec.
+3.2).
+Figure 5. Performance of selection strategies over training rounds on PolitiFact. The x-axis represents the training rounds, while the y-axis represents the detector’s AUC.
 Table 3 demonstrates that omitting any component leads to a decrease in performance for certain style conversions attack.
-While omitted due to space limitation, it can be confirmed that our sampling strategy converges faster than other sampling strategies (See Figure 8 in Appendix).
-(a) Diversity
-(b) Adversarialness
-(c) Coherency
-Figure 5: Qualitative analyses of AdStyle.
-(a) Histogram of diversity in embedding ($\mathbf{z}_{c}$) of conversion prompts sampled by our method and adversarial only selection.
-(b) Histogram of adversarialness scores ($s_{adv}^{c}$) for selected and unselected prompts (c) Histogram of coherency scores for selected and unselected prompts
-Qualitative analyses.
-First, we verified that our sampling strategy effectively selects diverse style-conversion prompts with high adversarialness and coherency.
-Figure 5(a) visualizes the diversity of prompts selected by our method compared to those chosen solely based on adversarialness scores (i.e., Adversarial-only Selection, the third model in our ablation study).
-We measure diversity using the average cosine similarity of every pair of selected prompts’ embeddings, $\mathbf{z}_{c}$ (Eq.
-2):
-$\text{Diveristy}(\mathcal{C})$ $=1-{1\over|\mathcal{C}|}\sum_{(c_{i},c_{j})\in\mathcal{C}}{\text{sim}(\mathbf{z}_{c_{i}},\mathbf{z}_{c_{j}})},$ (3)
-where $\mathcal{C}$ is the set of prompt pairs, $\text{sim}(\cdot)$ represents the cosine similarity.
-The result in the Figure 5(a) indicate that our strategy results in a more diverse set of augmentations compared to selection based on adversarialness alone.
-Figure 5(b) and 5(c) illustrate the Adversarialness and Coherency, respectively, of our selected prompts compared to remaining unselected prompts.
-Adversarialness was measured using the $s_{adv}^{c}$ score (Section 3.3), and coherency was calculated as the cosine similarity between the original text and its augmented version using semantic BERT embeddings Chanchani and Huang (2023).
-We can also observe that, for both metrics, prompts selected through AdStyle exhibit higher values compared to unselected prompts.
-This suggests that our strategy effectively selects prompts by considering both adversarialness and coherency.
-Next, we qualitatively analyzed the characteristics of our style-conversion prompts.
-Figure 6 shows examples of prompts generated and selected by AdStyle.
-Previous work relied on prompts that were more conventional, such as “neutral" or “sensational." However, our prompts included creative phrases beyond typical human suggestions, which proved to be more adversarial to the detector.
-Currently, we use a fixed format for the initial set of prompts, which results in the generation of prompts with a similar format.
-We expect that using a more diverse initial set of prompts will allow the LLM to optimize prompts within a wider search space.
-Round 0 P1: Rewrite the following article in a nonsensical and absurdly exaggerated tone with a hint of horror P2: Rewrite the following article in a sarcastic and mocking tone P3: Rewrite the following article in a chaotic and disorganized tone Round 1 P1: Rewrite the following article in a haunting and macabre tone with a sense of impending horror and madness P2: Rewrite the following article in a cryptic and enigmatic tone P3: Rewrite the following article in a malevolent and apocalyptic tone with a sense of impending doom and destruction, while also incorporating elements of surrealism and existential dread
-Figure 6: Examples of our adversarial style-conversion prompts selected for the PolitiFact dataset
-5 Conclusion
-This paper presents a robust fake news detection method that withstands various style-conversion and paraphrasing attacks through adversarial style-conversion.
-Unlike traditional detectors that use predefined, agnostic augmentations, AdStyle employs tailored augmentations that shift samples in the direction of the detector’s current decision boundary using style-conversion prompts, functioning similarly to adversarial noise.
-Among the various prompt candidates generated by the LLM, we selected an efficient set of prompts for training by considering diversity, coherency, and adversarialness.
-As a result, we were able to train a detector that exhibits high robustness and generalizability against a wide range of attacks.
-We believe that our work helps filter fake news and contribute to a better exchange of information in online.
-Appendix A APPENDIX
-A.1 Data & Implementation Details
-Dataset overview.
-This paper utilizes a diverse range of datasets sourced from publicly available fake news detection resources, selected for their unique topical focuses.
-We utilize PolitiFact and Gossipcop, drawn from the FakeNewsNet benchmark Shu et al.
-(2020), addressing political claims and celebrity rumors respectively.
-Furthermore, we include Constraint Felber (2021), a dataset focused on COVID-19 related social media posts.
-Each dataset is randomly divided into an 80% training set and a 20% test set.
-Detailed statistics for these datasets can be found in Table 3.
-Table 3: Statistics of fake news datasets.
-Dataset PolitiFact GossipCop Constraint
-# of News Articles 774 7,916 8,418
-# of Real News 399 3,958 4,406
-# of Fake News 375 3,958 4,012
-Implementation details.
-For consistency, experiments replicating existing baselines maintained fixed settings for learning, backbone network, and other relevant parameters.
-Augmented variations for UDA were generated using back-translation through German.
-RADAR employed adversarial training to learn paraphrases, utilizing the maximization of the binary cross-entropy loss of a fake news detector as a reward signal, with the T5-large model serving as the paraphraser.
-Named entities for ENDEF were extracted using the bert-base-NER model from Hugging Face.
-For SheepDog, augmentations were generated following the original paper’s methodology, using the prompt format illustrated in Figure 7 and four tones: "objective and professional," "neutral," "emotionally triggering," and "sensational".
-We also utilized the same four tones and prompt format as our predefined initial prompt set.
-Two V100 GPUs were utilized for all experiments.
-Rewrite the following article in a/an [tone]: [news article]
-Figure 7: Prompt for style-conversion. The “tone" part will be filled with the desired tone, and the “news article" part will contain the original news text.
-A.2 Further Results on Evaluation.
-Result on selection strategies across rounds.
-Figure 8 shows the AUC for each selection strategy over different training rounds.
-Random Selection leads to slow convergence and unstable training, while Adversarial-only Selection, limited by its narrow focus on adversarial prompts, can negatively impact performance in later stages of training due to a lack of diversity in the generated prompts.
-Notably, without score trajectories, the model struggles to generate prompts that contribute effectively to the training process.
-Figure 8: Performance of selection strategies over training rounds on the PolitiFact dataset. The x-axis represents the training rounds, while the y-axis represents the detector’s AUC.
-Table 4: Performance comparison of different attack scenarios on the PolitiFact Dataset. (A: Adversarial prompt, B: Summarization prompt, C: In-Context prompt D: Adversarial Paraphraser)
+Notably, excluding the score trajectory component when selecting the style-conversion prompt proved to be the most detrimental to performance.
+This finding suggests that the LLM can effectively identify conversion prompts that may confuse the detector through the style-conversion prompt and confusion score pairs as in-context demonstrations.
+Moreover, selectively choosing conversion prompts identified by the LLM based on specific criteria resulted in further performance improvements.
+Our sampling strategy also facilitates faster convergence compared to alternative sampling methods (See Figure 5).
+Table 4. Performance comparison of different LLM-based baselines on the PolitiFact Dataset. (NY: The New York Times, TS: The Sun, NE: National Enquirer)
+Model CNN NY TS NE Clean
+GPT-3.5 0.5820 0.6242 0.6173 0.5274 0.7037
+GPT-4 0.7133 0.6748 0.7025 0.6021 0.7823
+4.3. Performance Analysis
+We here conduct analysis on how AdStyle demonstrates robust and high performance across various scenarios and how it effectively enhances the detector’s performance.
+Comparison with LLM-based baselines.
+AdStyle enhanced the detector’s performance by leveraging the reasoning abilities of advanced large language models like GPT-3.5.
+To determine whether the capabilities of an advanced large language model alone are sufficient for the fake news detection task, we compared AdStyle with other LLM-based baselines: zero-shot and in-context learning-based inference with GPT-3.5.
+In the case of the in-context learning baseline, one example each of fake news and real news was provided as in-context demonstrations.
+The example instruction prompt for the LLM-based baselines are as follows:
+Table 4 presents the comparison results on the PolitiFact dataset.
+Using LLMs with prompting alone (Figure 6) makes it challenging to accurately determine the authenticity of news.
+Instead of ‘directly’ leveraging the LLM’s text generation capability for inference, it is more effective to use it as an augmentation tool to provide additional training signals, as demonstrated by our model.
+Does the following contain real or fake news? Answer in one word with either ‘Real’or ‘Fake’: [news article]
+Figure 6. Instruction prompt for LLM-based baselines. The ”news article” section contains the original news text.
+Table 5. Comparison under attack scenarios with Gemini-Pro on the PolitiFact dataset.
+Model CNN NY TS NE Average
+Vanilla 0.7913 0.8039 0.8882 0.8397 0.8308
+UDA 0.7216 0.7863 0.8657 0.8119 0.7964
+RADAR 0.8023 0.7805 0.8764 0.8423 0.8254
+ENDEF 0.7730 0.7537 0.8439 0.8058 0.7941
+SheepDog 0.8487 0.8926 0.9174 0.8998 0.8896
+Ours 0.8821 0.9120 0.9295 0.9241 0.9120
+Robustness against a other LLM backbones.
+We assessed the robustness by examining its performance against style conversion attacks when the attacker utilizes a different LLM backbone, such as Gemini-Pro (Reid et al., 2024).
+As shown in Table 5, AdStyle consistently outperforms the baselines even with a different LLM backbone, suggesting that the proposed approach is not overly reliant on recognizing the specific style of content generated by the backbone.
+Robustness against other possible attack scenarios.
+To validate the model’s robustness against various attacks, we have conducted comparison experiments with diverse attack scenarios to deceive the detector by altering textual style of inputs: (1) Adversarial prompt: Given a news article and its label, the prompt instructs the LLM to rewrite the article to evade detection as the given label.
+(2) Summarization prompt: A prompt instructing the LLM to summarize the news article without incorporating stylistic guidance.
+(3) In-Context prompt: A prompt providing an example of a recent, real CNN article, instructing the LLM to rewrite the given article in the same style.
+(4) Adversarial Paraphraser: The attack utilizes a paraphraser adversarially trained on the training dataset.
+The example prompt for each attack is described in Figures 12 to 14 in Appendix.
+Based on the results in Table 6, we confirm that our proposed model still demonstrates a performance improvement compared to other baselines against these attacks.
+Table 6. Comparison under attack scenarios on the PolitiFact Dataset (A: Adversarial prompt, B: Summarization prompt, C: In-Context prompt D: Adversarial Paraphraser).
 Model A B C D
@@
 Ours 0.9456 0.9416 0.9425 0.9212
-Comparison on other possible attack scenarios.
-We have conducted additional comparison experiments with more diverse attack scenarios: (1) Adversarial prompt: Given a news article and its label, the prompt instructs the LLM to rewrite the article to evade detection as the given label.
-(2) Summarization prompt: A prompt instructing the LLM to summarize the news article without incorporating stylistic elements.
+(a) Diversity
+(b) Adversarialness
+(c) Coherency
+Figure 7.
+Qualitative analyses with y-axis representing (a) diversity in embedding ($\mathbf{z}_{c}$) of conversion prompts sampled by AdStyle and adversarial-only selection; (b) Adversarialness ($s_{adv}^{c}$) for selected and unselected prompts; and (c) Coherency for selected and unselected prompts.
+Figure 8. Comparison of augmented samples generated by our model and SheepDog: (A) Histogram of lexical diversity, (B) Histogram of perplexity, and (C) Perplexity across rounds.
+Effectiveness of the selection strategy.
+We here empirically verified that our selection strategy (Sec.
+3.3) effectively selects diverse style-conversion prompts with high adversarialness and coherency.
+Figure 7(a) visualizes the diversity of prompts selected by our method compared to those chosen solely based on adversarialness scores (i.e., Adversarial-only Selection, the third model in our ablation study).
+We measure diversity using the average cosine similarity of every pair of selected prompts’ embeddings, $\mathbf{z}_{c}$ (Eq.
+2):
+(Eq. 4) $\text{Diveristy}(\mathcal{C})$ $=1-{1\over|\mathcal{C}|}\sum_{(c_{i},c_{j})\in\mathcal{C}}{\text{sim}(\mathbf{z}_{c_{i}},\mathbf{z}_{c_{j}})},$
+where $\mathcal{C}$ is the set of prompt pairs, $\text{sim}(\cdot)$ represents the cosine similarity.
+The result in the Figure 7(a) indicate that our strategy results in a more diverse set of augmentations compared to selection based on adversarialness alone.
+Round 0
+P1: Rewrite the following article in a nonsensical and absurdly exaggerated tone with a hint of horror
+P2: Rewrite the following article in a sarcastic and mocking tone
+P3: Rewrite the following article in a chaotic and disorganized tone
+Round 1
+P1: Rewrite the following article in a haunting and macabre tone with a sense of impending horror and madness
+P2: Rewrite the following article in a cryptic and enigmatic tone
+P3: Rewrite the following article in a malevolent and apocalyptic tone with a sense of impending doom and destruction, while also incorporating elements of surrealism and existential dread
+Figure 9. Example adversarial style-conversion prompts selected for the PolitiFact dataset.
+Figure 7(b) and 7(c) illustrate the Adversarialness and Coherency, respectively, of our selected prompts compared to remaining unselected prompts.
+Adversarialness was measured using the $s_{adv}^{c}$ score (Section 3.3), and coherency was calculated as the cosine similarity between the original text and its augmented version using semantic BERT embeddings (Chanchani and Huang, 2023).
+We can also observe that, for both metrics, prompts selected through AdStyle exhibit higher values compared to unselected prompts.
+This suggests that our strategy effectively selects prompts by considering both adversarialness and coherency.
+Analysis on style-conversion prompts.
+Finally, we conducted a qualitative analysis of our style-conversion prompts to understand what characteristics of the augmented samples contribute to improving the detector’s robustness.
+Interestingly, when we compare the augmented samples from two models: AdStyle and SheepDog, we found that our model produced sentences with significantly higher perplexity according to the language model backbone used by the detector than SheepDog model (9.05 vs.
+4.26, see Figure 8B), even though ours have lower lexical diversity (0.503 vs.
+0.634, see Figure 8A).
+In other words, our generated samples featured sentence structures that the detector’s language model likely had not encountered during pretraining.
+By providing the detector with inputs that have a variety of sentence structures and styles it has not previously seen, the detector naturally becomes robust against style-conversion attacks.
+In addition, when we measured the changes in perplexity of the augmented samples across training rounds (see Figure 8C), the perplexity increased over rounds and eventually converged at a certain point.
+This indicates that iterative exploration over multiple rounds is effective in creating augmentations that increasingly challenge the detector.
+Figure 9 shows examples of prompts generated and selected by AdStyle.
+In contrast, our prompts included creative phrases beyond typical human-crafted suggestions, making them adversarial to the detector.
+The round 0 result shows a prompt asking for “(P1) a nonsensical and absurdly exaggerated tone with a hint of horror”, while round 1 offers a lengthier prompt mentioning, for example, “(P3) a malevolent and apocalyptic tone with a sense of impending doom and destruction, while also incorporating elements of surrealism and existential dread.” Previous work relied on conventional prompts, such as “neutral” or “sensational.” The LLM is particularly well-suited for exploring such a vast array of candidates.
+Currently, we use a fixed format for the initial set of prompts, which results in the generation of prompts with a similar format.
+We expect that using a broader set of prompts will allow the LLM to optimize prompts within a wider search space.
+5. Conclusion
+We presented a robust fake news detection method that effectively withstands paraphrasing attacks through adversarial style conversion.
+Unlike traditional detectors that use predefined and agnostic augmentations, AdStyle employs tailored augmentations that shift samples in the direction of the detector’s current decision boundary using style-conversion prompts, functioning similarly to adversarial noise.
+Among the various candidates for the LLM prompts, we selected an efficient set of prompts for training by considering diversity, coherence, and adversarialness.
+Consequently, we were able to train a detector that exhibits high robustness and generalizability against a wide range of attacks.
+This robustness ensures that the detector can effectively identify and mitigate various forms of style-conversion attacks, regardless of the target news outlet.
+We believe that our research, combined with the shared codes, contributes to better filtering of fake news online.
+We aim to support the broader goal of ensuring the integrity of information in the digital age.
+Ethical Consideration
+We recognize several ethical concerns in this work.
+First, our method may inherit biases from the training data, which can be mitigated by carefully selecting style-conversion results that preserve textual coherence.
+Second, while our approach aims to detect fake news, similar techniques could be misused to enhance misinformation—an ongoing challenge in adversarial research.
+Lastly, our detection methods might influence news production by constraining creative writing styles, highlighting the need for further research on the balance between algorithmic moderation and information systems.
+Acknowledgments
+J.G.
+Lee and M.
+Cha are co-corresponding authors.
+We extend our gratitude to Fangzhao Wu, Wenchao Dong, and the anonymous reviewers for their insightful feedback on our work.
+This research was supported by the National Research Foundation of Korea (NRF) grant (RS-2022-00165347).
+Appendix A APPENDIX
+A.1. Baseline Implementation
+For consistency, experiments replicating existing baselines maintained fixed settings for learning, backbone network, and other relevant parameters.
+Augmented variations for UDA were generated using back-translation through German.
+RADAR employed adversarial training to learn paraphrases, utilizing the maximization of the binary cross-entropy loss of a fake news detector as a reward signal, with the T5-large model serving as the paraphraser.
+Named entities for ENDEF were extracted using the bert-base-NER model from Hugging Face.
+For SheepDog, augmentations were generated following the original paper’s methodology, using the prompt format illustrated in Figure 10 and four tones: ”objective and professional,” ”neutral,” ”emotionally triggering,” and ”sensational”.
+Rewrite the following article in a/an [tone]: [news article]
+Figure 10. Prompt for style-conversion. The “tone” part will be filled with the desired tone, and the “news article” part will contain the original news text.
+A.2. Examples of Style-Conversion Prompt
+Figure 11 showcases examples generated by the selected style-conversion prompt, demonstrating how fake news content can be modified while preserving its meaning.
+These examples demonstrate the impact of style transformation on fake news detection.
+Original: Updated: ASPEN 2014 Senator John McCain is meeting with his top campaign advisers and donors here…
+Rewrite the following article in a sarcastic and mocking tone: Oh, how exciting! Senator John McCain is gracing us with his presence in Aspen, meeting with his elite group of advisers and donors…
+Rewrite the following article in a chaotic and disorganized tone: Senator John McCain is like, in Aspen, meeting with his campaign peeps and rich donors, you know?…
+Rewrite the following article in a nonsensical and absurdly humorous tone with a touch of surrealism and a sprinkle of silliness : Senator John McCain found himself in a pickle.
+He was surrounded by his top campaign advisers and donors in the mystical town of Aspen…
+Figure 11. Generated example input created using the selected style-conversion prompt.
+A.3. Prompt Examples of Other Attack Scenarios
+We have conducted comparison experiments with more diverse attack scenarios: (1) Adversarial prompt: Given a news article and its label, the prompt instructs the LLM to rewrite the article to evade detection as the given label.
+(2) Summarization prompt: A prompt instructing the LLM to summarize the news article without incorporating stylistic guidance.
 (3) In-Context prompt: A prompt providing an example of a recent, real CNN article, instructing the LLM to rewrite the given article in the same style.
+Here we provide the example prompt for each attack prompt as figures.
 (4) Adversarial Paraphraser: The attack utilizes a paraphraser adversarially trained on the training dataset.
-The formats for each attack prompt are described in Figures 9 to 11.
-Based on the experimental results in Table 4, we confirm that our proposed model demonstrates a non-trivial performance improvement compared to other baselines against these attacks.
-<EXAMPLES>: [news article] Determine whether the given <EXAMPLES> content is fact or fake, mentioning the style of the article that led to that decision following format: Answer: {Fact or Fake} Reason: {one-sentence reason }
-Your current article is: [news article] But this article is detected as [Answer] due to the following [Reason].
+¡EXAMPLES¿: [news article]
+Determine whether the given ¡EXAMPLES¿ content is fact or fake, mentioning the style of the article that led to that decision following format:
+Answer: {Fact or Fake}
+Reason: {one-sentence reason}
+Your current article is: [news article]
+But this article is detected as [Answer] due to the following [Reason].
 Based on the above information, rewrite a new improved article not to be detected as [news label], maintaining the original content, as follows:
-Figure 9: The format of Adversarial prompt.
-First, use the prompt above to extract the answer and reason of factuality of the given news article.
-Then, use the following prompt for paraphrasing to avoid detection based on the identified reason.
-The "news article" section contains the original news text, and the "news label" section contains the corresponding news label.
+Figure 12.
+Example input for Adversarial prompt model.
+First, use the prompt above to extract the answer and reason for the authenticity of the given news.
+Next, utilize the prompt below to rephrase the news to evade detection based on the given reason.
+The ”news article” section contains the original news text, and the ”news label” part contains the corresponding label.
 Summarize the following article, ensuring the content remains the same: [news article]
-Figure 10: The format for Summarization prompt. The “news article" part will be filled with the original news text.
-<EXAMPLES>: [news article example] Rewrite the following article as the writing style of <EXAMPLES> : [news article]
-Figure 11: The format for In-Context prompt. The "news article example" section is filled with an example from a specific publisher, while the "news article" section contains the original news text.
-Comparison with LLM-based Baselines.
-We compared AdStyle with other LLM-based baselines: GPT-3.5 OpenAI (2022), GPT-4 OpenAI (2023), and In-Context Learning Brown et al.
-(2020).
-OpenAI APIs are utilized to predict news article labels, with the temperature set to 0 to ensure stable and consistent predictions.
-In the case of In-context Learning, one example each of fake news and real news was provided.
-The prompt formats for the LLM-based baselines are as follows:
-Does the following contain real or fake news? Answer in one word with either ‘Real’or ‘Fake’: [news article]
-Figure 12: Prompt for LLM-based baselines. The "news article" section contains the original news text.
-Table 5: Performance comparison of different LLM-based baselines on the PolitiFact Dataset. (NY: The New York Times, TS: The Sun, NE: National Enquirer)
-Model CNN NY TS NE Clean
-GPT-3.5 0.5820 0.6242 0.6173 0.5274 0.7037
-GPT-4 0.7133 0.6748 0.7025 0.6021 0.7823
-In-Context 0.6954 0.6504 0.6875 0.5754 0.7383
-Ours 0.9174 0.9444 0.9520 0.9460 0.9460
-Full result on style-conversion attack.
-Table 6 reports the AUC for each style-conversion attack.
-Paired sample t-tests were conducted to compare our method with the second-best baseline, Sheepdog, across different ratios for each dataset.
-Except for the 25% ratio in the Gossipcop dataset, all p-values were significantly smaller than 0.05, indicating statistically significant differences.
-Table 6: Performance comparison with AdStyle in four different style—conversion attacks.
-For the attack scenario, we report the average AUC of four style-conversion attacks.
-The best results are marked in bold.
-Our model demonstrates a significant performance improvement over all text-based fake news detectors in diverse style-conversion attack scenarios.
-(NY: The New York Times, TS: The Sun, NE: National Enquirer )
-CNN Politifact Gossipcop Constraint
-0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
-Vanilla 0.5664 0.6566 0.8127 0.6841 0.7913 0.8518 0.8563 0.8175 0.8056
-UDA 0.5906 0.7174 0.8009 0.7294 0.7859 0.8599 0.8305 0.8594 0.9021
-RADAR 0.7205 0.7301 0.7926 0.7683 0.8138 0.8666 0.8099 0.8400 0.8873
-ENDEF 0.5426 0.6679 0.7588 0.7288 0.7839 0.8624 0.7695 0.8666 0.8353
-SheepDog 0.6305 0.8269 0.8683 0.7472 0.8393 0.8706 0.9200 0.9022 0.9544
-AdStyle 0.7547 0.8822 0.9174 0.8195 0.8429 0.8744 0.9344 0.9430 0.9646
-NY Politifact Gossipcop Constraint
-0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
-Vanilla 0.5899 0.7334 0.8687 0.6750 0.7797 0.8465 0.8522 0.8939 0.8868
-UDA 0.6189 0.7644 0.8744 0.7207 0.7795 0.8558 0.8745 0.9184 0.9297
-RADAR 0.6966 0.7853 0.8586 0.7481 0.7960 0.8574 0.8701 0.8974 0.9176
-ENDEF 0.6152 0.7630 0.8150 0.7227 0.7837 0.8609 0.8341 0.9067 0.8918
-SheepDog 0.6310 0.7990 0.9090 0.7420 0.8309 0.8650 0.8937 0.9254 0.9645
-AdStyle 0.7599 0.8940 0.9444 0.8118 0.8389 0.8686 0.9243 0.9594 0.9720
-TS Politifact Gossipcop Constraint
-0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
-Vanilla 0.6913 0.6859 0.8789 0.7068 0.7918 0.8430 0.7822 0.8727 0.8974
-UDA 0.6646 0.8072 0.8883 0.7488 0.7904 0.8613 0.8662 0.9145 0.9446
-RADAR 0.7271 0.7173 0.8909 0.7594 0.8062 0.8620 0.8564 0.8882 0.9132
-ENDEF 0.7001 0.8132 0.8444 0.7603 0.7909 0.8639 0.8445 0.8955 0.9043
-SheepDog 0.7070 0.8408 0.9096 0.7580 0.8395 0.8688 0.8962 0.9217 0.9665
-AdStyle 0.8444 0.9002 0.9520 0.8117 0.8347 0.8751 0.9219 0.9521 0.9727
-NE Politifact Gossipcop Constraint
-0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
-Vanilla 0.6432 0.6856 0.8591 0.7020 0.7875 0.8399 0.7833 0.8853 0.9067
-UDA 0.6557 0.7892 0.8622 0.7450 0.7902 0.8595 0.8747 0.9191 0.9423
-RADAR 0.7428 0.7269 0.8864 0.7573 0.8026 0.8556 0.8775 0.8919 0.9161
-ENDEF 0.6924 0.7877 0.8355 0.7502 0.7897 0.8589 0.8456 0.9112 0.9025
-SheepDog 0.6636 0.8268 0.9166 0.7521 0.8331 0.8632 0.8606 0.9257 0.9667
-AdStyle 0.8029 0.8913 0.9460 0.8166 0.8391 0.8702 0.9092 0.9579 0.9771
+Figure 13. Example input for Summarization prompt model. The “news article” part is filled with the original news text.
+¡EXAMPLES¿: [news article example] Rewrite the following article as the writing style of ¡EXAMPLES¿ : [news article]
+Figure 14.
+Example input for In-Context prompt model.
+The ”news article example” section is filled with an example from a specific publisher, while the ”news article” section contains the original news text.